### v1

# Ground Every Sentence: Improving Retrieval-Augmented LLMs with Interleaved Reference-Claim Generation

Abstract
Retrieval-Augmented Generation (RAG) has been widely adopted to enhance Large Language Models (LLMs) in knowledge-intensive tasks. Recently, Attributed Text Generation (ATG) has attracted growing attention, which provides citations to support the model’s responses in RAG, so as to enhance the credibility of LLM-generated content and facilitate verification. Prior methods mainly adopt coarse-grained attributions, linking to passage-level references or providing paragraph-level citations. However, these methods still fall short in verifiability and require certain time costs for fact checking. This paper proposes a fine-grained ATG method called ReClaim (Refer & Claim), which alternates the generation of references and answers step by step. Unlike traditional coarse-grained attribution, ReClaim allows the model to add sentence-level fine-grained citations to each answer sentence in long-form question-answering tasks. Our experiments encompass various training and inference methods and multiple LLMs, verifying the effectiveness of our approach.
1 Introduction
Retrieval-Augmented Generation (RAG) Lewis et al. (2020) is a technique that integrates information retrieval with natural language generation to enhance the performance and accuracy of large language model (LLMs) responses. However, the RAG system still encounters challenges related to verifiability and credibility. To address these issues, attributed text generation (ATG) Bohnet et al. (2022) has been proposed. ATG aims to improve RAG systems in terms of: 1) Credibility, as explicit citations can reduce hallucinations; 2) Verifiability, making it easier for users to verify the answer.
Previous efforts on ATG mainly focus on passage-level (Thoppilan et al., 2022) or paragraph-level references (Menick et al., 2022; Nakano et al., 2021; Gao et al., 2023b). Although these attribution methods have contributed to improving the verifiability and credibility of model responses, current methods often focus on relatively coarse-grained attributions, which may contain a significant amount of irrelevant information. This increases the time cost for fact-checking.
Figure 1: The task setup for ReClaim. Given a question and reference passages from a large retrieval corpus. The model then generates a text response with fine-grained citations.
In this paper, we propose ReClaim, which generates attributed text with interleaving references and answers for RAG systems, as is shown in Figure 1. This method enables sentence-level fine-grained attributions for the model’s answer in long-form question-answering using the RAG system.
To improve the model’s text generation with citations, we constructed a training dataset, fine-tuned the model, and employed strategies for alternating citation and answer generation. We also added decoding constraints to prevent inconsistencies between citations and source reference passages.
Results of experimental results indicate that our method not only matches but also surpasses existing baseline in some metrics, particularly improving the attribution quality, which refers to the extent to which citations support the answer text. Additionally, our method reduces the length of citations, thereby reducing the effort needed for fact-checking.
Our contributions are summarized as follows:
  1. 1.
    We propose a method called ReClaim, which alternately generates citations and answer sentences, to enable large models to generate answer with citations.
  2. 2.
    For ReClaim, we constructed a training dataset and fine-tuned the model using different approaches to improve its attribution capability.
  3. 3.
    Through multiple experiments, we demonstrated the effectiveness of our method in enhancing the model’s verifiability and credibility.
2 Related Work
Retrieval-Augmented Generation
In this paper, we use the RAG (Retrieval-Augmented Generation) system to generate answer with citations. The RAG system was proposed to combine information retrieval with generation models for tasks such as question answering and knowledge generation. Similarly, this system has been widely applied to handle complex tasks that require extracting information from a large number of documents, including open-domain question answering, dialogue systems, and information summarization Izacard and Grave (2021); Karpukhin et al. (2020).
Long-form Text Question Answering
Our work primarily focuses on the long-form question answering (LFQA) task within the RAG system. Unlike short-form QA Rajpurkar et al. (2016); Joshi et al. (2017), which concentrates on extracting concise facts, LFQA aims to generate comprehensive answers that require a deep understanding of the context and the integration of information from multiple sources. A notable work in this field is the ELI5 (Explain Like I’m Five) dataset Fan et al. (2019), which challenges models to provide straightforward and comprehensible explanations for complex questions. Similarly, the ASQA (Answer Summaries for Questions which are Ambiguous) dataset Stelmakh et al. (2022) requires models to generate abstractive summaries from multiple answer passages, ensuring that the synthesized answer is both coherent and informative.
Generate Text with Citation
Many current works have proposed various methods to generate answer text with citations. These methods differ in their approaches to attribution and the granularity of the citations.
LaMDA Thoppilan et al. (2022) provides attribution for the entire response in the form of URLs pointing to entire documents. WebGPT Nakano et al. (2021) and GopherCite Menick et al. (2022) use reinforcement learning from human preferences to enable LLMs to answer questions while providing snippet citations. ALCE Gao et al. (2023b) goes further by providing one or more input documents as attribution for each sentence in the answer, in a manner similar to cross-referencing.
In addition to the aforementioned methods that add citations directly during answer generation, there are some methods that focus on finding citations afterward. RARR Gao et al. (2023a) uses two steps, Research and Revision, to add attribution to any text post hoc and to edit the original answer based on the found citations.
3 Method
This study aims to generate text with fine-grained citations by integrating citations with answers in the form of Chain-of-Thought (CoT) Wei et al. (2022). We introduce our approach ReClaim in Section 3.1, and the following sections detail the specific implementation of ReClaim.
3.1 ReClaim: Interleaving Reference and Claim
Our task can be formally expressed as follows: Given a query q and several reference passages retrieved by the RAG system $\mathcal{D}$, the LLMs is required to generate an output $\mathcal{O}$ = { $\textit{r}_{\textit{1}}$, $\textit{c}_{\textit{1}}$, $\textit{r}_{\textit{2}}$, $\textit{c}_{\textit{2}}$$\textit{r}_{\textit{n}}$, $\textit{c}_{\textit{n}}$, where $\mathcal{O}$ consists of n sentence-level fine-grained references $\textit{r}_{\textit{1}}$, …, $\textit{r}_{\textit{n}}$, which represent the fine-grained citations coming from reference passages and n claims $\textit{c}_{\textit{1}}$, …, $\textit{c}_{\textit{n}}$, which are portions of the model’s response generated based on these references. Each reference $\textit{r}_{\textit{i}}$ corresponds to and serves as substantiation for the corresponding claim $\textit{c}_{\textit{i}}$. All the claims together form the model’s complete answer to the question.
In our work, references are surrounded and denoted by the tags <reference> and </reference>, while claims are demarcated using the tags <claim> and </claim>. Then, they are connected in the form of CoT to produce attributed text response.
During generation, the model needs to alternately generate the reference and claim parts. In the experimental process, it was found that the model encounters several issues when generating references and claims: 1) The references generated by the model are not always completely consistent with the context of the retrieved reference passages; 2) The claims generated by the model do not always attribute well to the corresponding references. Therefore, in the following sections, we will study how to improve the generation of references and claims.
3.2 Training Dataset Construction
To improve the model’s ability to generate the references and corresponding claims, we constructed a specialized fine-tuning dataset. This dataset was built based on the WebGLM-QA Liu et al. (2023) dataset which consists of 43579 high-quality data samples for the train split. These data samples contain rich reference passages and detailed long-form answer to the question. We performed dataset construction steps and data cleansing processes on this dataset, ultimately resulting in 9433 training samples suitable for our task format.
We adopted a two-stage method to construct our training dataset: 1) Use ChatGPT (gpt-3.5-turbo) OpenAI (2022) to automatically segment the answers in the WebGLM-QA dataset, and identified the corresponding citations in the reference passages based on the segmented clauses; 2) Constructed the attributed answers using the clauses and their corresponding citations in the form of CoT as described in Section 3.1.
To ensure the quality of attribution in the constructed answers, we refined the preliminary training dataset: 1) Segmented the corresponding citation of each clause and performed string matching in the reference passages to filter out mismatched data items where the citation did not align with the original reference passages; 2) Used a natural language inference (NLI) Honovich et al. (2022) model to judge the entailment relationship between the citations (as the premise) and the clauses (as the hypothesis). If there is no implication relationship, it indicates that the citation does not contain sufficient information to support the corresponding clause. Therefore, we need to filter out such data items to ensure a relatively high level of attribution quality.
Through these steps, we ultimately constructed a fine-tuned dataset that ensures text consistency of the references and high attribution quality, providing a foundation for subsequent research.
Samples Average Length
Answer Citation Passages
9433 93.6 141.8 282.8
Table 1: Statistics of the training dataset.
3.3 Unified Generation
[…]
3.4 Interleaving Generation
During the claim generation stage, since the model has already selected sufficiently granular reference text to follow, which contain the answer information, the full input context is not required. Therefore, the ReClaim w/IG (IG stands for Interleaving Generation) method trains separate models for the generation of the reference parts and the claim parts, and alternates between the two models during answer generation, adjusting the input to each model accordingly. The whole generation process is shown in Figure 2. Specifically, we adopted the following steps to train the models and generate the answer:
Figure 2: The generation process of ReClaim w/IG. Based on the given questions and the reference passages retrieved, the model alternately generates the reference parts and the claim parts in a step-by-step manner. For these two stages of generation, distinct datasets are constructed to train the models, which alternately switches between the models and the input context during inference.
Reference Generation
During the generation of the reference parts in the output, the model needs to generate the next reference based on the complete input context and previous output. We define this generation process as $ReferGen=\{\textit{r}_{\textit{i+1}}\ |\ Prompt,\{\textit{r}_{\textit{1}},\textit{c}_{\textit{1}},...,\textit{r}_{\textit{i}},\textit{c}_{\textit{i}}\}\}$, where $\textit{r}_{\textit{i+1}}$ refers to the reference part generated in the next stage, $Prompt$ refers to the complete input context containing instructions, query and reference passages, and $\textit{r}_{\textit{1}},\textit{c}_{\textit{1}},...,\textit{r}_{\textit{i}},\textit{c}_{\textit{i}}$ refer to the previously generated references and claims. As the training of the model for generating the reference parts does not require masking parts of the input context information, we follow the same approach as in the Section 3.3, using the training data constructed in Section 3.2 to fine-tune the model.
Claim Generation
During the generation of the claim parts in the output, the model only needs to generate the next claim based on the previous output. We define this generation process as $ClaimGen=\{\textit{c}_{\textit{i+1}}\ |\{\textit{r}_{\textit{1}},\textit{c}_{\textit{1}},...,\textit{r}_{\textit{i}},\textit{c}_{\textit{i}},\textit{r}_{\textit{i+1}}\}\}$, where $\textit{c}_{\textit{i+1}}$ refers to the claim part generated in the next stage, and $\textit{r}_{\textit{1}},\textit{c}_{\textit{1}},...,\textit{r}_{\textit{i}},\textit{c}_{\textit{i}},\textit{r}_{\textit{i+1}}$ refer to the previously generated references and claims. We use the training data constructed in Section 3.2 and split it into the data format that conforms to our described generation process ClaimGen. This formatted dataset is then used to fine-tune the model.
After generating the two models mentioned above, we can alternate between these models during the inference stage while simultaneously adjusting the model input according to the defined generation process. This allows us to separately generate the reference and claim parts, ultimately producing attributed text output.
3.5 Decoding Constraints
To ensure the generated reference parts align with the source reference passages, we impose decoding constraints through the following three steps:
Sentence Segmentation and Encoding
We segment the reference passages into individual sentences. Then, we use the model tokenizer to encode these sentences into vectors. Each vector representation of a sentence serves as the smallest unit for generating the reference parts.
Constructing Prefix Tree
The encoded vectors are transformed into a list format and organized into a Prefix tree Fredkin (1960) structure. Employing such a structure to store our reference sentences facilitates the choice of the next token in subsequent generation steps.
Constrained Inference
During the model inference stage for generating reference parts, we select the token with the highest generation probability that satisfies the current prefix tree path as the next output token. This process continues until reaching a leaf node. Upon reaching a leaf node, the model either select another prefix tree path for output or begin the claim generation. This approach allows us to select a complete and consistent sentence from the reference passages as part of our current reference each time.
4 Experimental Protocol
In this section, we employ the GPT models and several open-source models to validate the effectiveness of our method across multiple evaluation dimensions. We conduct a comprehensive analysis by assessing the performance of our approach on various metrics.
4.1 Evaluation Datasets
Our method primarily targets long-form question answering task within the RAG system, aiming to generate attributed answer. Therefore, we selected ASQA Stelmakh et al. (2022) and ELI5 Fan et al. (2019) as our main benchmark datasets. By testing on these datasets, we can comprehensively assess the effectiveness of our method in generating accurate, coherent, and well-attributed long-form answers.
4.2 Evaluation Metrics
Building on our previous task definition, we focus on evaluating the model-generated outputs in three key areas:
Answer Quality
We first evaluate the quality of the answers generated by the model. We concatenate all claim parts in order to form the answer to the question, and follow the ALCE evaluation method to calculate the correctness and fluency of the answers. For answer correctness, in the ASQA dataset, we use Exact Match Recall (EM Rec.) to measure the percentage of golden short answers contained in the generated answers. In the ELI5 dataset, we adopt Claim Recall (Claim Rec.) to measure the percentage of key claims included in the answers. To evaluate answer fluency, we use MAUVE Pillutla et al. (2021) to measures the similarity between output and gold answer.
Citation Quality
Similar to ALCE, we employ the AutoAIS Bohnet et al. (2022) to measure the citation quality. Our citation quality is also measured by two metrics: 1) Correct Attribution Score (CAS), which determines if the answers is entirely supported by cited sentences; 2) Citation Redundancy Score (CRS), which identifies any redundant citation sentences.
For CAS, We use the TRUE Honovich et al. (2022) model to compute the entailment relationship between each reference part and the corresponding claim part. The final CAS score is the proportion of sentences predicted as correctly attributed among all the sentences in the answer. For CRS, since our method allows the model to select multiple contextual reference sentences for the same claim, we need to determine if the reference contains redundant sentences. The final CRS score is the proportion of non-redundant sentences relative to all sentences in the references.
Verifiability and Credibility
We employ two metrics to measure the Verifiability: 1) Citation Length, where shorter citation text typically reduces the time needed for fact-checking; 2) Attribution Ratio (AR), which represents the proportion of sentences in the output that are attributed.
We use Consistency Ratio (CR) to measure the Credibility, which determines the text consistency between the reference parts and the reference passages through string matching. Higher CR usually means the model generates fewer hallucinations when outputting the reference parts.
Method Model ASQA ELI5 Average
Fluency Correct. Citation Fluency Correct. Citation Fluency Correct. Citation
MAUVE EM Rec. CAS CRS MAUVE Claim Rec. CAS CRS MAUVE Claim Rec. CAS CRS
ALCE VANILLA 66.8 40.4 73.6 72.5 57.2 12.0 51.1 50.0 62.0 26.2 62.4 61.3
SUMM 70.0 43.3 68.8 61.8 40.2 12.5 51.5 48.2 55.1 27.9 60.2 55.0
SNIPPET 69.8 41.4 65.3 57.4 62.9 14.3 50.4 45.0 66.4 27.9 57.9 51.2
ORACLE 64.4 48.9 74.5 72.7 59.4 21.3 57.8 56.0 61.9 35.1 66.2 64.4
0-shot GPT-4o 72.9 52.8 74.8 51.6 37.3 19.9 63.5 30.7 55.1 36.4 69.2 41.2
3-shot Vicuna-7B 81.4 50.5 74.0 70.0 38.6 13.9 73.2 59.7 60.0 32.2 73.6 64.9
Vicuna-13B 85.0 48.5 76.1 68.2 35.5 15.6 74.3 57.4 60.3 32.1 75.2 62.8
Llama3-8B 90.1 50.7 77.7 62.1 61.3 17.9 78.3 45.3 75.7 34.3 78.0 53.7
ChatGPT 74.9 52.6 72.5 63.4 27.8 17.8 68.6 50.8 51.4 35.2 70.6 57.1
GPT-4o 91.3 56.6 77.4 58.0 29.7 21.1 70.2 36.8 60.5 38.9 73.8 47.4
[…]
Vicuna-13B 75.9 46.3 57.7 51.1 64.1 19.0 52.5 37.6 70.0 32.7 55.1 44.4
Llama3-8B 88.6 50.4 65.4 57.4 73.7 19.4 59.3 43.8 81.2 34.9 62.4 50.6
ReClaim w/IG Vicuna-13B 88.7 49.1 84.4 74.7 66.8 16.9 77.4 54.6 77.8 33.0 80.9 64.7
Llama3-8B 86.4 50.5 87.7 77.1 66.8 16.8 84.3 61.8 76.6 33.7 86.0 69.5
Table 2: Results on ALCE benchmark Gao et al. (2023b) and ReClaim. Definitions for Fluency, Correct. and Citation are in Section 4.2.
4.3 Methods and Baseline
To evaluate the effectiveness of the proposed method, we tested several models using various generation approaches. For each experimental setup, we used the ChatGPT results from the ALCE method as our baseline. Additionally, for the two test datasets, we adopted the Oracle-5 paragraphs provided by the ALCE as the reference passages in our model input.
Prompting
We directly prompt the model to generate answer text with citations. For GPT-4o OpenAI (2023), we employ both zero-shot and three-shot prompting Brown et al. (2020), while for GPT-3.5-turbo and open-source models such as Llama3-8B-Instruct Touvron et al. (2023) and the vicuna Chiang et al. (2023) series, we use 3-shot prompting. This approach allows us to comprehensively assess the performance and effectiveness of various models under different prompting scenarios.
Fine-tuned Model
[…]
As outlined in Section 3.4, we fine-tune the same model, generating two separate models: ReferModel and ClaimModel. We completed our training on four 80GB A800 GPUs. For the ReferModel, we performed full fine-tuning with a learning rate of 2e-5 over 3 epochs. For the ClaimModel, we employed Lora tuning Hu et al. (2021) with a learning rate of 5e-5 over 5 epochs. During the inference phase, we alternated between these two models to interleavingly generate the reference and claim parts. To prevent the model from generating overly short or long answers, we set constraints on the number of reference-claim pairs. For the ASQA dataset, the model must generate between 2 to 5 pairs. For the ELI5 dataset, due to its complexity, the model must generate between 4 to 6 pairs.
To investigate the roles of the two fine-tuned models, ReferModel and ClaimModel, in the ReClaim w/IG method, we conducted ablation experiments using the Llama3-8B-Instruct model. In these experiments, we alternately used only the fine-tuned ReferModel or ClaimModel, while employing the base model with 3-shot prompting to generate the other model’s output.
For these two fine-tuning methods, we adopted the constrained decoding approach described in Section 3.5 to limit the model’s output in the reference parts.
Method ASQA ELI5 Average
Fluency Correct. Citation Fluency Correct. Citation Fluency Correct. Citation Length
MAUVE EM Rec. CAS CRS MAUVE Claim Rec. CAS CRS MAUVE Claim Rec. CAS CRS
ALCE 64.4 48.9 74.5 72.7 59.4 21.3 57.8 56.0 61.9 35.1 66.2 64.4 -
Citation-only 83.5 53.8 65.1 54.6 47.4 17.4 75.1 52.4 65.5 35.6 70.1 53.5 78.9
Claim-only 43.6 57.7 87.8 72.0 52.7 17.8 81.4 52.2 48.2 37.8 84.6 62.1 121.7
ReClaim w/IG 86.4 50.5 87.7 77.1 66.8 16.8 84.3 61.8 76.6 33.7 86.0 69.5 86.8
Table 3: The impact of different components under the ReClaim w/IG method on model generation results. Citation-only indicates using the fine-tuned ReferModel to generate references and the base model to generate claims. Conversely, Claim-only indicates using the fine-tuned ClaimModel to generate claims and the base model to generate references.
Method ASQA ELI5
Length Consistency Attri. Length Consistency Attri.
Citation Claim CR AR Citation Claim CR AR
ALCE 536.3 85.5 100.0 91.3 660.0 98.09 100.0 96.9
[…]
Table 4: The generated text length, consistency of references, and proportion of attributed answer sentences in different methods. CD indicates the use of constrained decoding, while UD indicates the absence of constrained decoding.
5 Experiment Results
In the experiments, we wish to answer two research questions: $RQ1)$ How to improve the quality of answers and citations? $RQ2)$ Can ReClaim enhance the verifiability and credibility of RAG-based question answering?
5.1 How to Improve the Quality of Answers and Citations?
The overall performance is presented in Table 2:
Prompting Works
Experimental results show that direct prompt models can achieve satisfactory outcomes. Compared to the ALCE method using Oracle-5 reference passages, our approach often matches or improves answer flunecy, correctness and citation correctness (CAS).
In all our prompting experiments, GPT-4o achieved the highest average correctness score (38.9) under 3-shot prompting. Additionally, Llama3-8B-Instruct achieved satisfactory results. Although it did not surpass the GPT model in terms of average correctness score, it demonstrated the best performance in fluency and CAS. Compared to the ALCE method, our model maintained response correctness while improving average fluency and CAS by 22.3% and 17.8%, respectively.
Although our method performs worse in CRS, the finer granularity of our citations minimizes the impact of redundant content. In contrast to paragraph-level citations, where each redundant citation introduces an entire paragraph of irrelevant content, our redundant citations only introduce a single sentence of irrelevant content. This does not significantly increase the fact-checking cost. Moreover, in many cases, redundant but continuous text helps in locating the citation in the original reference passages and aids in its understanding.
[…]
ReClaim w/IG Improves Attribution
Experimental results indicate that the ReClaim w/IG method outperforms other methods in two citation quality metrics while maintaining high fluency and correctness scores. Specifically, when using the Llama3-8B-Instruct model, our method outperforms the ALCE method with ChatGPT by 23.7%, 30.0%, and 7.9% in fluency, CAS, and CRS metrics, respectively, and only shows a 4.0% decrease in correctness.
[…]
As shown in Table 3, the experimental results demonstrate that the ReClaim w/IG method achieves the best performance in metrics of fluency and citation quality when compared to fine-tuning only the RefModel or the ClaimModel methods. This indicates the effectiveness of our fine-tuned dual-model approach. While the Claim-only method, which fine-tunes only the ClaimModel, yields the highest average answer correctness score (37.8), it comes at the cost of increased response length and reduced fluency. Additionally, numerous experiments indicate that improved citation quality often leads to a slight decrease in answer correctness. This is expected, as better attribution tends to constrain the source of information for model responses to the selected citations.
5.2 Can ReClaim Enhance the Verifiability and Credibility of RAG-based Question Answering?
To evaluate the verifiability of model responses and the impact of decoding constraints on their credibility, we measured the average length of citations, the consistency of citation texts (CR), and the proportion of sentences with attribution (AR) under different methods when using constrained decoding (CD) and unconstrained decoding (UD). The experimental results are shown in Table 4.
ReClaim Improves Verifiability
Experimental results indicate that ReClaim’s average citations length is only about 20% of the length produced by the ALCE method, significantly reducing the fact-checking time cost. Compared to the ALCE method, our method ensures specific attribution for each response sentence, further improving the verifiability of model’s answer.
Decoding Constraints Improve Credibility
[…]
6 Conclusion
We propose a attributed text generation method, ReClaim, which adds sentence-level fine-grained citations to model-generated answer in RAG systems. This approach alternates between generating citations and answer sentences through Prompting or by fine-tuning the model.
The results show that our method improves citation quality while maintaining answer quality compared to the baseline method. Additionally, our approach significantly reduces the length of citations, thus decreasing the time cost required for fact-checking and further enhancing the verifiability of the model’s responses. Moreover, by using constrained decoding during citation generation, we ensure that each citation is composed of exact sentences from the source passages.
Although this paper focuses on long-form question answering task, we hope that our method can be extended to other tasks such as short-form question answering and Multi-document Summarization to improve the verifiability and credibility of model-generated responses.
Limitations
In this paper, our training dataset was exclusively targeted at long-form question-answering task, which reduces the generalization ability of our fine-tuning methods. Additionally, the construction of our training dataset did not account for the influence of irrelevant passages in the context on the model’s ability to find citations. This may weaken the model’s capability to filter out irrelevant information.
While our approach allows the model to synthesize information from multiple reference sentences for attribution, we did not specifically enhance this capability in our training data construction process and the model’s inference process.
Ethics Statement
We hereby acknowledge that all authors of this work are aware of the provided ACL Code of Ethics and honor the code of conduct.
Datasets Source
All original datasets used for training and testing were sourced from open and publicly accessible resources, and they are all approved for use in research purposes, thereby minimizing the risk of sensitive information leakage. While we employed LLMs for automated processing during the construction of training dataset, data cleansing was performed to prevent the introduction of additional noise and bias. We solely utilize the constructed dataset for model training. Although paragraph retrieval is not the focus of our work, information retrieved from large corpora may introduce some noise and bias into the responses generated by LLMs. To address these issues, we will optimize the data construction process and explore methods for retrieving noise-free and unbiased information.
AI assistants
AI assistants (ChatGPT) were solely used to improve the grammatical structure of the text.
Appendix A Notation Table
Definition
Task Formulation
$\mathcal{q}$ The given question.
$\mathcal{D}$ A set of reference passages, where $d\in\mathcal{D}$ is a passage.
$\mathcal{O}$ The response generated based on the question $\mathcal{q}$ and passages $\mathcal{D}$, composed of $\mathcal{O}=\{\textit{r}_{\textit{1}},\textit{c}_{\textit{1}},...,\textit{r}_{\textit{i}},\textit{c}_{\textit{i}}\}$
$r$ A portion of the citation, comprising certain sentences from $\mathcal{D}$.
$c$ A portion of the answer, formed by concatenating all $c$ in sequence as a response to $\mathcal{q}$.
Methods
ALCE A benchmark for Automatic LLMs’ Citation Evaluation.
ReClaim-Unified End-to-end data training and one-step generation.
ReClaim w/IG Independent model training and Interleaving Generation.
Citation-only Use the fine-tuned model to generate references and the base model with 3-shot prompting to generate claims.
Claim-only Use the base model with 3-shot prompting to generate references and the fine-tuned model to generate claims.
UD Use of constrained decoding.
CD Absence of constrained decoding.
Metrics
MAUVE Measuring the gap between neural text and human text using divergence frontiers.
EM Rec. Exact match recall rate of gold short answers in the text generated by the model.
Claim Rec. Recall rate of generated claims in the text generated by the model.
CAS Correct attribution score, the proportion of sentences predicted as correctly attributed among all the sentences in the answer.
CRS Citation redundancy score, the proportion of non-redundant sentences relative to all sentences in the references.
CR Consistency ratio, the text consistency between the reference parts and the reference passages through string matching.
AR Attribution ratio, the proportion of sentences in the output that are attributed.
Table 5: A notation table.
In Table 5, we list the notations and abbreviations in this paper, along with their definitions.
Appendix B Details of Training Dataset Construction
B.1 Citation Selection
During the training data structuring phase, we initially employ ChatGPT to automatically segment the answers within the WebGLM-QA dataset into clauses. Subsequently, for each of these clauses, we undertake a search for corresponding sentence-level, fine-grained citations from the provided reference passages. The prompt we use is as follows:
Upon obtaining answer clauses and their respective citations, structure the response in a cot format, as illustrated in Table 6.
B.2 Citation Filtering
After identifying the citation sentences corresponding to each answer clause, we need to filter out citation texts that diverge from the original reference passages or lack sufficient information to substantiate the answer clauses, the algorithm we use is as Algorithm 1.
Algorithm 1 Citation Filtering
1:Data: Question $\mathcal{q}$, Reference passages $\mathcal{D}$, Answer $\mathcal{a}=\{(r_{i},c_{i})~{}|~{}r_{i}\in R,c_{i}\in C\}$, NLI model $M$
2:for each $Tuple(\mathcal{q},\mathcal{D},\mathcal{a})$ do
3:     $flag\leftarrow 1$
4:     for each citation $r_{i}$ and its corresponding clause $c_{i}$ do
5:         for $sentence\in~{}r_{i}$ do
6:              if sentence not in $\mathcal{D}$ then
7:                  $flag\leftarrow 0$
8:              end if
9:         end for
10:         $m\leftarrow M(r_{i},c_{i})$
11:         if $m=0$ then
12:              $flag\leftarrow 0$
13:         end if
14:     end for
15:     if $flag=0$ then
16:         remove $Tuple(\mathcal{q},\mathcal{D},\mathcal{a})$
17:     end if
18:end for
Question Why does a dryer sheet help with my staticy shirt?
References [1] Dryer sheets are specifically designed to help reduce static cling in clothes by neutralizing the electric charges that build up during the drying process. Just toss a sheet in with your clothes, and they should come out nice and static-free when they’re done drying.
[2] So, if you if you want your clothes to be ready to wear, but they’re always covered in pet hair or are super clingy, then you definitely need to prevent static cling! Don’t worry, we’ve all been there. Bounce Dryer Sheets will help your clothes lead a no-strings-attached lifestyle in no time. Just toss a sheet into the dryer with your clothes, and leave that static cling behind!
[3] However, this coating can stick to your dryer and can be bad for the environment as well. Those looking to know how to reduce static on clothes without dryer sheets can find some solid, eco-friendly dryer sheet alternatives in many different places.
[4] The most common way people know how to prevent dryer static on clothes is with dryer sheets. Dryer sheets are sheets that are coated in a fabric softener full of positively charged electrons. These bond to the negatively charged ones and keep static from happening.
[5] Dryer static on clothes is one of the most annoying parts of doing laundry. It might seem like an unavoidable problem, but we have good news! There are some simple steps you can take to keep your laundry from becoming overly clingy or giving you static shocks.
Answer According to the citation: <reference> The most common way people know how to prevent dryer static on clothes is with dryer sheets. Dryer sheets are sheets that are coated in a fabric softener full of positively charged electrons. These bond to the negatively charged ones and keep static from happening. </reference> We can know that: <claim> Dryer sheets are coated in a fabric softener full of positively charged electrons, which bond to the negatively charged ones in your clothes and keep static from happening. </claim>
Table 6: Example of the training dataset.
Appendix C Experiment Settings
C.1 Details of Datasets
The statistic of the test datasets of ReClaimis shown in Table 7.
ASQA
ASQA Stelmakh et al. (2022) is the first long-form question answering dataset that focuses on ambiguous factoid questions (see Table 9). It contains 4,353 samples for the train split and 948 samples for the dev split. For ALCE Gao et al. (2023b), ASQA questions were paired with passages from Wikipedia passages (2018-12-20 snapshot) which purportedly contained the answers.
ELI5
ELI5 Fan et al. (2019) dataset is a benchmark in natural language processing designed for long-form question answering tasks, focusing on complex and explanatory questions (see Table 9). It comprises 270K threads from the Reddit forum “Explain Like I’m Five” (ELI5) where an online community provides answers to questions which are comprehensible by five year olds. For ALCE Gao et al. (2023b), ELI5 questions were paired with passages from Sphere Piktus et al. (2021), a filtered version of Common Crawl.
Dataset Samples Question Type #passages Average Length
Question Passages Answer
ASQA 948 Factoid (ambiguous) Wikipedia (21M) 9.0 500.0 71.8
ELi5 1000 Why/How/What Sphere (899M) 16.5 500.0 121.5
Table 7: Statistics of the test datasets.
Input According to the citation: <reference> The most common way people know how to prevent dryer static on clothes is with dryer sheets. Dryer sheets are sheets that are coated in a fabric softener full of positively charged electrons. These bond to the negatively charged ones and keep static from happening. </reference>
Output We can know that: <claim> Dryer sheets are coated in a fabric softener full of positively charged electrons, which bond to the negatively charged ones in your clothes and keep static from happening. </claim>
Table 8: Example of the training dataset for claim generation.
ASQA Question: Who has the highest goals in world football?
Reference Passages: [1] "Football Player of the Century", by IFFHS International Federation of Football History and Statistics, 1999, "South America Football Player of the Century", by IFFHS International Federation of Football History and Statistics. Pelé’s 1281 goals are recognized by FIFA as the highest total achieved by a professional footballer, although the Soccer Statistic Foundation (rssf) recognizes only 767 goals in official mode, occupying the third place after Josef Bican (805) and Romario (772). For his part, Maradona has been named the best soccer player in World Cup history both by The Times and FourFourTwo, publication that also rewarded him as the "Best
[2] for Christmas, but died less than 2 weeks before that, at the age of 88. Josef Bican Josef "Pepi" Bican (25 September 1913 —— 12 December 2001) was a Czech-Austrian professional footballer who played as a striker. Rec.Sport.Soccer Statistics Foundation (RSSSF) estimates that he scored at least 805 goals in all competitive matches, which would make him the most prolific scorer of all time. Having scored a record 600 league goals and at least 1468 goals overall, the International Federation of Football History & Statistics (IFFHS) awarded Bican the "Golden Ball" as the greatest goalscorer of the last century. He
[3] …
Answer: The players with the highest all-time goals and highest men’s and women’s international football goals differ. The player with the highest all-time men’s football goals is Josef Bican, who in 2020 was recognized by FIFA, the international governing body of football, as the record scorer with an estimated 805 goals. Christine Sinclair has the highest goals in women’s international football with 187 and is the all-time leader for international goals scored for men or women. Cristiano Ronaldo and Ali Daei are currently tied for leading goalscorer in the history of men’s international football with 109.
ELI5 Question: How are firms like snapchat, uber etc valued so highly while still not making a profit? Do venture capitalists not expect some form of repayment within a number of years?
Reference Passages: [1] reporting that the Los Angeles-based company is aiming to raise \$500 million at a valuation of \$16 billion to \$19 billion, making it the third most highly valued tech start-up backed by venture capitalists. The Chinese handset maker Xiaomi is valued at \$45 billion, while Uber is estimated to be valued at about \$40 billion, according to data from CB Insights. Read MoreVC investment hits \$86B thanks to Uber, Xiaomi Snapchat was valued at \$10 billion in August, according to a Dow Jones report. Some of its investors from previous rounds include Benchmark, Lightspeed Venture Partners and Kleiner Perkins Caufield
[2] Ever wondered how highly valued technology giants like Google and Facebook were able to grow so fast and pay their employees so well in such a short amount of time, or how still growing start-ups like Uber are able to lose 1.2 billion US dollars in just the first half of this year alone and still command a valuation upwards of 50 billion US dollars? The answer lies with a special category of investment activity known as venture capital. Venture capitalists are professional investors who invest in a number of highly scalable high-risk technology ventures hoping to make a multi-fold
[3] …
Answer: Yes. Did you watch The Social Network? They went a while before introducing ads, so they could make money, as they needed to establish their brand and amass users. Once you have dedicated users, introducing ads won’t deter most, but if you are still new, having ads will deter a lot. The same goes for Uber, it’s not that they aren’t making money, it’s that they are reinvesting a ton of it to make their service better.
Table 9: Example of the test datasets.
C.2 Details of Methods
Prompting
In prompting setting, given a question and reference passages, we simply prompt the model to generate the attributed answer in a prescribed format. The prompt we use is as follows:
[…]
ReClaim w/IG
In ReClaim w/IG setting, we train two separate models for generating the reference and claim parts of attributed answers. Specifically, we train the model using the constructed training dataset in Appendix B as ReferModel for producing reference parts. For ClaimModel, which generates claim parts, we construct a new form of training data based on the original dataset. The detailed structure of this tailored training data is shown in Table 8. In Table 10, we provide an illustration of an interleaving generation instance.
Step Model Input Output
1 ReferModel Instrction: Given the Question and References below, provide an answer for the Question that is generated using information exclusively from the References(some may be irrelevant). …
# Question: Who has the highest goals in world football?
# References: "Football Player of the Century", by IFFHS International Federation of Football History and Statistics, 1999, "South America Football Player of the Century", by IFFHS International Federation of Football History and Statistics. Pelé’s 1281 goals are recognized by FIFA as the highest total achieved by a professional footballer, although the Soccer Statistic Foundation (rssf) recognizes only 767 goals in official mode, occupying the third place after Josef Bican (805) and Romario (772). …
Output:
According to the citation: <reference> Pelé’s 1281 goals are recognized by FIFA as the highest total achieved by a professional footballer, although the Soccer Statistic Foundation (rssf) recognizes only 767 goals in official mode, occupying the third place after Josef Bican (805) and Romario (772). </reference>
2 ClaimModel According to the citation: <reference> Pelé’s 1281 goals are recognized by FIFA as the highest total achieved by a professional footballer, although the Soccer Statistic Foundation (rssf) recognizes only 767 goals in official mode, occupying the third place after Josef Bican (805) and Romario (772). </reference> We can know that: <claim> Pelé’s 1281 goals are recognized by FIFA as the highest total achieved by a professional footballer, although the Soccer Statistic Foundation (rssf) recognizes only 767 goals in official mode, occupying the third place after Josef Bican (805) and Romario (772). </claim>
3 ReferModel Instrction: Given the Question and References below, provide an answer for the Question that is generated using information exclusively from the References(some may be irrelevant). …
# Question: Who has the highest goals in world football?
# References: "Football Player of the Century", …
Output: According to the citation: <reference> Pelé’s 1281 goals are recognized by FIFA … </reference> We can know that: <claim> Pelé’s 1281 goals are recognized by FIFA … </claim>
According to the citation: <reference> Rec.Sport.Soccer Statistics Foundation (RSSSF) estimates that he scored at least 805 goals in all competitive matches, which would make him the most prolific scorer of all time. </reference>
4 ClaimModel According to the citation: <reference> Pelé’s 1281 goals are recognized by FIFA … </reference> We can know that: <claim> Pelé’s 1281 goals are recognized by FIFA … </claim> According to the citation: <reference> Rec.Sport.Soccer Statistics Foundation (RSSSF) estimates that he scored at least 805 goals in all competitive matches, which would make him the most prolific scorer of all time. </reference> We can know that: <claim> Rec.Sport.Soccer Statistics Foundation (RSSSF) estimates that he scored at least 805 goals in all competitive matches, which would make him the most prolific scorer of all time. </claim>
Table 10: Example of the interleaving generation.

### v2

@@
 Retrieval-Augmented Generation (RAG) has been widely adopted to enhance Large Language Models (LLMs) in knowledge-intensive tasks.
-Recently, Attributed Text Generation (ATG) has attracted growing attention, which provides citations to support the model’s responses in RAG, so as to enhance the credibility of LLM-generated content and facilitate verification.
-Prior methods mainly adopt coarse-grained attributions, linking to passage-level references or providing paragraph-level citations.
-However, these methods still fall short in verifiability and require certain time costs for fact checking.
-This paper proposes a fine-grained ATG method called ReClaim (Refer & Claim), which alternates the generation of references and answers step by step.
-Unlike traditional coarse-grained attribution, ReClaim allows the model to add sentence-level fine-grained citations to each answer sentence in long-form question-answering tasks.
-Our experiments encompass various training and inference methods and multiple LLMs, verifying the effectiveness of our approach.
+To enhance credibility and verifiability in RAG systems, Attributed Text Generation (ATG) is proposed, which provides citations to retrieval knowledge in LLM-generated responses.
+Prior methods mainly adopt coarse-grained attributions, with passage-level or paragraph-level references or citations, which fall short in verifiability.
+This paper proposes ReClaim (Refer & Claim), a fine-grained ATG method that alternates the generation of references and answers step by step.
+Different from previous coarse-grained attribution, ReClaim provides sentence-level citations in long-form question-answering tasks.
+With extensive experiments, we verify the effectiveness of ReClaim in extensive settings, achieving a citation accuracy rate of 90%.
+[^1] Code and datasets are public at: https://github.com/pdxthree/ReClaim
 1 Introduction
 Retrieval-Augmented Generation (RAG) Lewis et al.
-(2020) is a technique that integrates information retrieval with natural language generation to enhance the performance and accuracy of large language model (LLMs) responses.
+(2020) is a technique that integrates information retrieval with natural language generation to enhance the performance of large language model (LLMs) responses.
 However, the RAG system still encounters challenges related to verifiability and credibility.
@@
 Although these attribution methods have contributed to improving the verifiability and credibility of model responses, current methods often focus on relatively coarse-grained attributions, which may contain a significant amount of irrelevant information.
-This increases the time cost for fact-checking.
-Figure 1: The task setup for ReClaim. Given a question and reference passages from a large retrieval corpus. The model then generates a text response with fine-grained citations.
+This increases the time required for fact-checking.
+Figure 1: The task setup for ReClaim.
+Given question and reference passages from a large corpus.
+The LLM then generates a response with fine-grained citations.
+For detailed examples, see Table LABEL:table:testdatacase.
 In this paper, we propose ReClaim, which generates attributed text with interleaving references and answers for RAG systems, as is shown in Figure 1.
-This method enables sentence-level fine-grained attributions for the model’s answer in long-form question-answering using the RAG system.
-To improve the model’s text generation with citations, we constructed a training dataset, fine-tuned the model, and employed strategies for alternating citation and answer generation.
-We also added decoding constraints to prevent inconsistencies between citations and source reference passages.
-Results of experimental results indicate that our method not only matches but also surpasses existing baseline in some metrics, particularly improving the attribution quality, which refers to the extent to which citations support the answer text.
-Additionally, our method reduces the length of citations, thereby reducing the effort needed for fact-checking.
+This method enables sentence-level fine-grained attributions in long-form question-answering using the RAG system.
+To enhance the LLM’s performance in attributed text generation, we developed a training dataset and fine-tuned the LLM to facilitate sentence-level citation selection from given reference passages and subsequent answer generation.
+We implemented an alternating strategy between citation generation and answer text generation.
+We apply constrained decoding during LLM inference by encoding reference passages into a token-level prefix tree.
+This restricts the LLM to generate citations only along the tree’s paths, ensuring alignment with the reference passages and avoiding inconsistencies.
+The results of experiments demonstrate that ReClaim outperforms existing baselines.
+ReClaim significantly improves the citation quality, enabling the citations to better support the answer.
+Furthermore, ReClaim greatly reduces the verbosity of citations, thereby easing the fact-checking process.
 Our contributions are summarized as follows:
   1. 1.
-    We propose a method called ReClaim, which alternately generates citations and answer sentences, to enable large models to generate answer with citations.
+    We propose a method called ReClaim, which alternately generates citations and answer sentences, enabling LLM to produce answers with sentence-level citations, thus enhancing the LLM’s verifiability and credibility.
   2. 2.
-    For ReClaim, we constructed a training dataset and fine-tuned the model using different approaches to improve its attribution capability.
+    To enhance LLMs in sentence-level citation generation, we construct a dataset based on WebGLM-QA Liu et al.
+(2023) and ELI5 Fan et al.
+(2019) dataset.
+Then, we fine-tune Llama3-8B-Instruct Dubey et al.
+(2024) models for reference and claim generation respectively, achieving better citation quality compared to the baseline method with ChatGPT OpenAI (2022).
   3. 3.
-    Through multiple experiments, we demonstrated the effectiveness of our method in enhancing the model’s verifiability and credibility.
+    Through extensive experiments, we demonstrate the effectiveness of our method in enhancing the LLM’s verifiability and credibility, achieving performance comparable to much bigger models like ChatGPT.
 2 Related Work
@@
 (2016); Joshi et al.
-(2017), which concentrates on extracting concise facts, LFQA aims to generate comprehensive answers that require a deep understanding of the context and the integration of information from multiple sources.
-A notable work in this field is the ELI5 (Explain Like I’m Five) dataset Fan et al.
-(2019), which challenges models to provide straightforward and comprehensible explanations for complex questions.
-Similarly, the ASQA (Answer Summaries for Questions which are Ambiguous) dataset Stelmakh et al.
-(2022) requires models to generate abstractive summaries from multiple answer passages, ensuring that the synthesized answer is both coherent and informative.
-Generate Text with Citation
-Many current works have proposed various methods to generate answer text with citations. These methods differ in their approaches to attribution and the granularity of the citations.
+(2017), which concentrates on extracting concise facts, LFQA generates comprehensive answers that require deep contextual understanding and information integration from multiple sources Fan et al.
+(2019); Stelmakh et al.
+(2022); Malaviya et al.
+(2024).
+Attributed Text Generation
+Many current works propose various methods for generating answer text with citations, differing in their approaches to attribution and citation granularity.
 LaMDA Thoppilan et al.
@@
 (2023b) goes further by providing one or more input documents as attribution for each sentence in the answer, in a manner similar to cross-referencing.
-In addition to the aforementioned methods that add citations directly during answer generation, there are some methods that focus on finding citations afterward.
-RARR Gao et al.
-(2023a) uses two steps, Research and Revision, to add attribution to any text post hoc and to edit the original answer based on the found citations.
+Additionally, some work has focused on fine-tuning models to improve the generation of attributed answer text Huang et al.
+(2024a); Asai et al.
+(2023); Huang et al.
+(2024c).
+In addition to the aforementioned methods that add citations directly during answer generation, there are some works that focus on finding citations afterward Gao et al.
+(2023a).
+Some research further achieves better attribution performance through multiple retrievals and validations Li et al.
+(2024); Sun et al.
+(2023).
 3 Method
-This study aims to generate text with fine-grained citations by integrating citations with answers in the form of Chain-of-Thought (CoT) Wei et al.
+We introduce ReClaim, which aims to generate text with sentence-level citations.
+Specifically, ReClaim generates citations with answers alternatively, in a step-by-step manner.
+We first introduce the overview of ReClaim in Section 3.1, and then detail the specific implementation in Section 3.2 to 3.5.
+Figure 2: The generation process of ReClaim w/IG.
+Based on the given questions and the reference passages retrieved, the LLM alternately generates the reference parts and the claim parts in a step-by-step manner.
+For these two stages of generation, distinct datasets are constructed to train the base model, which alternately switches between the fine-tuned models and the input context during inference.
+3.1 ReClaim: Interleaving Reference and Claim
+Our task can be formally expressed as follows: Given a query q and several reference passages retrieved by the RAG system $\mathcal{D}$, the LLM is required to generate an output $\mathcal{O}$ = { $\textit{r}_{\textit{1}}$, $\textit{c}_{\textit{1}}$, $\textit{r}_{\textit{2}}$, $\textit{c}_{\textit{2}}$$\textit{r}_{\textit{n}}$, $\textit{c}_{\textit{n}}$, where $\mathcal{O}$ consists of n sentence-level fine-grained references $\textit{r}_{\textit{1}}$, …, $\textit{r}_{\textit{n}}$, which represent the fine-grained citations coming from reference passages (provided as text, not just source numbers) and n claims $\textit{c}_{\textit{1}}$, …, $\textit{c}_{\textit{n}}$, which are portions of the LLM’s response generated based on these references.
+Each reference $\textit{r}_{\textit{i}}$ serves as a substantiation of claim $\textit{c}_{\textit{i}}$, and together, all $\textit{c}_{\textit{i}}$ form the LLM’s complete answer to the question.
+During generation, the LLM alternates between generating references and claims.
+However, experiments revealed that the LLM faces several issues: 1) The generated references are not always consistent with the retrieved passages; 2) The generated claims do not always attribute well to the corresponding references.
+Therefore, in the following sections, we study how to improve the generation of references and claims.
+3.2 Training Dataset Construction
+To improve the LLM’s ability to choose the references and generate corresponding claims, we construct a specialized fine-tuning dataset based on the WebGLM-QA Liu et al.
+(2023) and ELI5 Fan et al.
+(2019) datasets.
+The training data input includes an instruction, a query, and reference passages.
+The output consists of multiple reference-claim pairs, with the reference being a sentence-level citation and the claim being a part of the answer.
+The reference and claim components are designed to train the LLM’s ability to select accurate and relevant fine-grained citations from the complete reference passages, and to generate fluent, coherent answer texts that are faithful to the selected citations.
+The steps of our dataset construction are detailed as follows:
+Reference Passages Retrieval
+In the WebGLM-QA Dataset, the reference passages are fine-grained texts closely aligned with the query.
+However, some research has shown that irrelevant reference texts significantly impact the quality of LLM-generated answers Wu et al.
 (2024).
-We introduce our approach ReClaim in Section 3.1, and the following sections detail the specific implementation of ReClaim.
-3.1 ReClaim: Interleaving Reference and Claim
-Our task can be formally expressed as follows: Given a query q and several reference passages retrieved by the RAG system $\mathcal{D}$, the LLMs is required to generate an output $\mathcal{O}$ = { $\textit{r}_{\textit{1}}$, $\textit{c}_{\textit{1}}$, $\textit{r}_{\textit{2}}$, $\textit{c}_{\textit{2}}$$\textit{r}_{\textit{n}}$, $\textit{c}_{\textit{n}}$, where $\mathcal{O}$ consists of n sentence-level fine-grained references $\textit{r}_{\textit{1}}$, …, $\textit{r}_{\textit{n}}$, which represent the fine-grained citations coming from reference passages and n claims $\textit{c}_{\textit{1}}$, …, $\textit{c}_{\textit{n}}$, which are portions of the model’s response generated based on these references.
-Each reference $\textit{r}_{\textit{i}}$ corresponds to and serves as substantiation for the corresponding claim $\textit{c}_{\textit{i}}$.
-All the claims together form the model’s complete answer to the question.
-In our work, references are surrounded and denoted by the tags <reference> and </reference>, while claims are demarcated using the tags <claim> and </claim>.
-Then, they are connected in the form of CoT to produce attributed text response.
-During generation, the model needs to alternately generate the reference and claim parts.
-In the experimental process, it was found that the model encounters several issues when generating references and claims: 1) The references generated by the model are not always completely consistent with the context of the retrieved reference passages; 2) The claims generated by the model do not always attribute well to the corresponding references.
-Therefore, in the following sections, we will study how to improve the generation of references and claims.
-3.2 Training Dataset Construction
-To improve the model’s ability to generate the references and corresponding claims, we constructed a specialized fine-tuning dataset.
-This dataset was built based on the WebGLM-QA Liu et al.
-(2023) dataset which consists of 43579 high-quality data samples for the train split.
-These data samples contain rich reference passages and detailed long-form answer to the question.
-We performed dataset construction steps and data cleansing processes on this dataset, ultimately resulting in 9433 training samples suitable for our task format.
-We adopted a two-stage method to construct our training dataset: 1) Use ChatGPT (gpt-3.5-turbo) OpenAI (2022) to automatically segment the answers in the WebGLM-QA dataset, and identified the corresponding citations in the reference passages based on the segmented clauses; 2) Constructed the attributed answers using the clauses and their corresponding citations in the form of CoT as described in Section 3.1.
-To ensure the quality of attribution in the constructed answers, we refined the preliminary training dataset: 1) Segmented the corresponding citation of each clause and performed string matching in the reference passages to filter out mismatched data items where the citation did not align with the original reference passages; 2) Used a natural language inference (NLI) Honovich et al.
-(2022) model to judge the entailment relationship between the citations (as the premise) and the clauses (as the hypothesis).
-If there is no implication relationship, it indicates that the citation does not contain sufficient information to support the corresponding clause.
-Therefore, we need to filter out such data items to ensure a relatively high level of attribution quality.
-Through these steps, we ultimately constructed a fine-tuned dataset that ensures text consistency of the references and high attribution quality, providing a foundation for subsequent research.
-Samples Average Length
+To enhance the LLM’s performance against irrelevant contexts, we sample a portion of the constructed WebGLM-QA training dataset, retrieve the top-100 passages using BM25, and calculate the similarity between the query and passages with a Reranker model Chen et al.
+(2024).
+Passages with high BM25 ranking but low re-rank score are selected as irrelevant text and added to the training data.
+For the ELI5 dataset, we use BM25 to retrieve the top-100 reference passages, which included some irrelevant noise to simulate real-world retrieval conditions.
+Additionally, we use the Reranker to re-rank the top 100 passages retrieved from a subset of the selected ELI5 dataset, producing a refined re-ranked dataset.
+Model Answer Generation
+For the WebGLM-QA dataset, we directly use the original model responses.
+For the ELI5 datasets, due to significant information discrepancies between the human answers and the retrieved passages, we provide the top-5 retrieved passages to the LLM (Llama-3.1-405B-Instruct Dubey et al.
+(2024)) to generate long-form answers.
+Multi-Stage Citation Search
+To select sentence-level fine-grained citations for the answers generated by the LLM, we employ a multi-stage citation search approach that moves from coarse to fine.
+We first enable the Llama-3.1-405B-Instruct model to automatically segment the text into clauses.
+For each clause, the model identifies the minimal set of citations from the reference passages that sufficiently support it.
+Then, we use NLI (Natural Language Inference) model Honovich et al.
+(2022) to evaluate the entailment relationship between the chosen citations (as the premise) and the corresponding clause (as the hypothesis).
+To enhance filtering precision, we set a threshold ($\theta=0.8$) to discard cases where the entailment probability between the chosen citation and the clause falls below the threshold.
+Dataset Samples Average Length
 Answer Citation Passages
-9433 93.6 141.8 282.8
-Table 1: Statistics of the training dataset.
+WebGLM-QA 4582 98.53 154.62 326.57
+WebGLM-QA Extend 2605 83.85 114.1 396.16
+ELI5 Default 3383 91.33 132.34 545.01
+ELI5 Rerank 2653 107.54 158.52 542.02
+Table 1: Statistics of the training dataset. For more details, please refer to the appendix B.
 3.3 Unified Generation
-[…]
+The ReClaim ${}_{\texttt{Unified}}$ method uses a simple fine-tuning and inference approach.
+It first performs instruction fine-tuning on the LLM using the dataset constructed in Section 3.2.
+Then, it uses the fine-tuned LLM to perform one-step generation.
+Based on the given query and reference passages, it directly outputs the attributed answer.
+This generation process can be described as: $UnifiedGen=\{\{\textit{r}_{\textit{1}},\textit{c}_{\textit{1}},...,\textit{r}_{\textit{i}},\textit{c}_{\textit{i}}\}\ |\ Query,Passages\}$.
 3.4 Interleaving Generation
-During the claim generation stage, since the model has already selected sufficiently granular reference text to follow, which contain the answer information, the full input context is not required.
-Therefore, the ReClaim w/IG (IG stands for Interleaving Generation) method trains separate models for the generation of the reference parts and the claim parts, and alternates between the two models during answer generation, adjusting the input to each model accordingly.
-The whole generation process is shown in Figure 2.
-Specifically, we adopted the following steps to train the models and generate the answer:
-Figure 2: The generation process of ReClaim w/IG.
-Based on the given questions and the reference passages retrieved, the model alternately generates the reference parts and the claim parts in a step-by-step manner.
-For these two stages of generation, distinct datasets are constructed to train the models, which alternately switches between the models and the input context during inference.
+During the claim generation stage, since the LLM has already selected sufficiently granular reference text to follow, which contain the answer information, the full input context is not required.
+Therefore, the ReClaim w/IG method trains separate LLMs for the generation of the reference parts and the claim parts, and alternates between the two LLMs during answer generation, adjusting the input to each LLM accordingly (IG represents the interleaving use of two fine-tuned LLMs for the iterative generation of references and claims).
+The entire generation process, as illustrated in Figure 2, involves the following steps to train the LLMs and generate the answer.
 Reference Generation
-During the generation of the reference parts in the output, the model needs to generate the next reference based on the complete input context and previous output.
+During the generation of reference parts, the LLM needs to generate the next reference based on the complete input context and previous output.
 We define this generation process as $ReferGen=\{\textit{r}_{\textit{i+1}}\ |\ Prompt,\{\textit{r}_{\textit{1}},\textit{c}_{\textit{1}},...,\textit{r}_{\textit{i}},\textit{c}_{\textit{i}}\}\}$, where $\textit{r}_{\textit{i+1}}$ refers to the reference part generated in the next stage, $Prompt$ refers to the complete input context containing instructions, query and reference passages, and $\textit{r}_{\textit{1}},\textit{c}_{\textit{1}},...,\textit{r}_{\textit{i}},\textit{c}_{\textit{i}}$ refer to the previously generated references and claims.
-As the training of the model for generating the reference parts does not require masking parts of the input context information, we follow the same approach as in the Section 3.3, using the training data constructed in Section 3.2 to fine-tune the model.
+As the training of the LLM for generating the reference parts does not require masking parts of the input context information, we follow the same approach as in the Section 3.3 to fine-tune the ReferModel.
 Claim Generation
-During the generation of the claim parts in the output, the model only needs to generate the next claim based on the previous output.
+During the generation of claim parts, the LLM only needs to generate the next claim based on the previous output.
 We define this generation process as $ClaimGen=\{\textit{c}_{\textit{i+1}}\ |\{\textit{r}_{\textit{1}},\textit{c}_{\textit{1}},...,\textit{r}_{\textit{i}},\textit{c}_{\textit{i}},\textit{r}_{\textit{i+1}}\}\}$, where $\textit{c}_{\textit{i+1}}$ refers to the claim part generated in the next stage, and $\textit{r}_{\textit{1}},\textit{c}_{\textit{1}},...,\textit{r}_{\textit{i}},\textit{c}_{\textit{i}},\textit{r}_{\textit{i+1}}$ refer to the previously generated references and claims.
-We use the training data constructed in Section 3.2 and split it into the data format that conforms to our described generation process ClaimGen.
-This formatted dataset is then used to fine-tune the model.
-After generating the two models mentioned above, we can alternate between these models during the inference stage while simultaneously adjusting the model input according to the defined generation process.
-This allows us to separately generate the reference and claim parts, ultimately producing attributed text output.
+We utilize the training dataset from Section 3.2, formatting it to align with our ClaimGen generation process, and then use this formatted dataset to fine-tune the ClaimModel.
 3.5 Decoding Constraints
-To ensure the generated reference parts align with the source reference passages, we impose decoding constraints through the following three steps:
+To ensure the generated reference parts align with the retrieved reference passages, we apply decoding constraints through the following three steps:
 Sentence Segmentation and Encoding
 We segment the reference passages into individual sentences.
-Then, we use the model tokenizer to encode these sentences into vectors.
+Then, we use the LLM tokenizer to encode these sentences into vectors.
 Each vector representation of a sentence serves as the smallest unit for generating the reference parts.
@@
 Constrained Inference
-During the model inference stage for generating reference parts, we select the token with the highest generation probability that satisfies the current prefix tree path as the next output token.
+During the inference stage for generating reference parts, we select the token with the highest generation probability that satisfies the current prefix tree path as the next output token.
 This process continues until reaching a leaf node.
-Upon reaching a leaf node, the model either select another prefix tree path for output or begin the claim generation.
+Upon reaching a leaf node, the LLM either select another prefix tree path for output or begin the claim generation.
 This approach allows us to select a complete and consistent sentence from the reference passages as part of our current reference each time.
 4 Experimental Protocol
-In this section, we employ the GPT models and several open-source models to validate the effectiveness of our method across multiple evaluation dimensions.
+In this section, we employ the GPT models and several open-source LLMs to validate the effectiveness of our method across multiple evaluation dimensions.
 We conduct a comprehensive analysis by assessing the performance of our approach on various metrics.
 4.1 Evaluation Datasets
-Our method primarily targets long-form question answering task within the RAG system, aiming to generate attributed answer.
-Therefore, we selected ASQA Stelmakh et al.
-(2022) and ELI5 Fan et al.
-(2019) as our main benchmark datasets.
-By testing on these datasets, we can comprehensively assess the effectiveness of our method in generating accurate, coherent, and well-attributed long-form answers.
+ASQA
+We evaluate the 948 samples from the ASQA dataset Stelmakh et al.
+(2022), selected by ALCE, and use the five oracle paragraphs provided by ALCE as reference passages, which are chosen from the top 100 retrieved passages representing the gold passages.
+ELI5
+We evaluate the 1,000 samples selected from ELI5 dataset Fan et al. (2019) by ALCE and use the five oracle passages as reference passages.
+EXPERTQA
+To test the LLM’s generalization ability under the standard RAG process, we select 1,000 samples from the EXPRTQA Malaviya et al.
+(2024) dataset.
+We follow the standard RAG procedure: using BM25 to retrieve the top-100 passages, then re-ranking them to select the top 5 passages.
 4.2 Evaluation Metrics
-Building on our previous task definition, we focus on evaluating the model-generated outputs in three key areas:
+Building on our previous task definition, we focus on evaluating the LLM-generated outputs in three key areas.
+Below, we introduce three evaluation dimensions, with more detailed information and calculation methods provided in Appendix C.3.
 Answer Quality
-We first evaluate the quality of the answers generated by the model.
 We concatenate all claim parts in order to form the answer to the question, and follow the ALCE evaluation method to calculate the correctness and fluency of the answers.
 For answer correctness, in the ASQA dataset, we use Exact Match Recall (EM Rec.) to measure the percentage of golden short answers contained in the generated answers.
-In the ELI5 dataset, we adopt Claim Recall (Claim Rec.) to measure the percentage of key claims included in the answers.
+In the ELI5 and EXPERTQA dataset, we adopt Claim Recall (Claim Rec.) to measure the percentage of key claims included in the answers.
 To evaluate answer fluency, we use MAUVE Pillutla et al.
-(2021) to measures the similarity between output and gold answer.
+(2021) to measure the similarity between output and gold answer.
 Citation Quality
@@
 (2022) to measure the citation quality.
-Our citation quality is also measured by two metrics: 1) Correct Attribution Score (CAS), which determines if the answers is entirely supported by cited sentences; 2) Citation Redundancy Score (CRS), which identifies any redundant citation sentences.
-For CAS, We use the TRUE Honovich et al.
-(2022) model to compute the entailment relationship between each reference part and the corresponding claim part.
-The final CAS score is the proportion of sentences predicted as correctly attributed among all the sentences in the answer.
-For CRS, since our method allows the model to select multiple contextual reference sentences for the same claim, we need to determine if the reference contains redundant sentences.
-The final CRS score is the proportion of non-redundant sentences relative to all sentences in the references.
-Verifiability and Credibility
-We employ two metrics to measure the Verifiability: 1) Citation Length, where shorter citation text typically reduces the time needed for fact-checking; 2) Attribution Ratio (AR), which represents the proportion of sentences in the output that are attributed.
-We use Consistency Ratio (CR) to measure the Credibility, which determines the text consistency between the reference parts and the reference passages through string matching.
-Higher CR usually means the model generates fewer hallucinations when outputting the reference parts.
-Method Model ASQA ELI5 Average
+Our citation quality is also measured by two metrics: 1) Correct Attribution Score (CAS), which determines whether the answer is entirely supported by cited sentences and is the most critical metric in our evaluation; 2) Citation Redundancy Score (CRS), which identifies any redundant citation sentences.
+We use the NLI Honovich et al.
+(2022) model to compute the entailment relationship between each reference part and the corresponding claim part, and the final CAS score is the proportion of correctly attributed sentences in the answer.
+We also use the NLI model need to determine if the reference contains redundant sentences, and the final CRS score is the proportion of non-redundant sentences relative to all sentences in the references.
+Verifiability
+We employ three metrics to measure the verifiability: 1) Citation length, where shorter citation text typically reduces the time needed for fact-checking; 2) Attribution Ratio (AR), which represents the proportion of sentences in the output that are attributed; 3) Consistency Ratio (CR), which represents the proportion of text consistency between the reference parts and the reference passages through string matching.
+Method Model ASQA ELI5 EXPERTQA
 Fluency Correct. Citation Fluency Correct. Citation Fluency Correct. Citation
 MAUVE EM Rec. CAS CRS MAUVE Claim Rec. CAS CRS MAUVE Claim Rec. CAS CRS
-ALCE VANILLA 66.8 40.4 73.6 72.5 57.2 12.0 51.1 50.0 62.0 26.2 62.4 61.3
-SUMM 70.0 43.3 68.8 61.8 40.2 12.5 51.5 48.2 55.1 27.9 60.2 55.0
-SNIPPET 69.8 41.4 65.3 57.4 62.9 14.3 50.4 45.0 66.4 27.9 57.9 51.2
-ORACLE 64.4 48.9 74.5 72.7 59.4 21.3 57.8 56.0 61.9 35.1 66.2 64.4
-0-shot GPT-4o 72.9 52.8 74.8 51.6 37.3 19.9 63.5 30.7 55.1 36.4 69.2 41.2
-3-shot Vicuna-7B 81.4 50.5 74.0 70.0 38.6 13.9 73.2 59.7 60.0 32.2 73.6 64.9
-Vicuna-13B 85.0 48.5 76.1 68.2 35.5 15.6 74.3 57.4 60.3 32.1 75.2 62.8
-Llama3-8B 90.1 50.7 77.7 62.1 61.3 17.9 78.3 45.3 75.7 34.3 78.0 53.7
-ChatGPT 74.9 52.6 72.5 63.4 27.8 17.8 68.6 50.8 51.4 35.2 70.6 57.1
-GPT-4o 91.3 56.6 77.4 58.0 29.7 21.1 70.2 36.8 60.5 38.9 73.8 47.4
+Prompting-Based
+ALCE ChatGPT 64.4 48.9 74.5 72.7 59.4 21.3 57.8 56.0 53.7 20.5 66.7 64.9
+Llama3-8B 79.2 55.2 54.7 54.6 45.4 20.5 42.8 39.3 56.7 20.2 51.4 47.8
+w/Rerank 81.3 55.3 79.1 76.4 37.6 20.0 59.5 53.8 45.3 19.8 68.7 60.2
+Post-hoc
+ClosedBook Llama3-8B 50.5 28.9 13.7 13.7 66.8 16.8 17.4 17.4 37.6 25.4 11.2 11.2
+Open-book Llama3-8B 73.7 53.2 52.1 52.1 26.6 20.8 31.9 32.0 58.7 21.0 36.9 36.8
+Training-Based
+Self-RAG Llama2-7B 70.6 38.7 53.3 66.2 33.1 9.7 23.1 33.9 19.0 12.5 9.4 12.1
+RS+RL Llama2-7B 84.4 47.7 75.5 69.4 43.6 19.1 59.1 51.6 46.6 15.3 67.8 60.1
+FRONT Llama2-7B 72.5 56.5 72.2 66.0 49.7 18.1 64.0 59.1 56.5 14.7 73.6 68.9
+Our Methods
+0-shot GPT-4o 72.9 52.8 74.8 51.6 37.3 19.9 63.5 30.7 59.4 18.5 73.4 26.5
 […]
-Vicuna-13B 75.9 46.3 57.7 51.1 64.1 19.0 52.5 37.6 70.0 32.7 55.1 44.4
-Llama3-8B 88.6 50.4 65.4 57.4 73.7 19.4 59.3 43.8 81.2 34.9 62.4 50.6
-ReClaim w/IG Vicuna-13B 88.7 49.1 84.4 74.7 66.8 16.9 77.4 54.6 77.8 33.0 80.9 64.7
-Llama3-8B 86.4 50.5 87.7 77.1 66.8 16.8 84.3 61.8 76.6 33.7 86.0 69.5
-Table 2: Results on ALCE benchmark Gao et al. (2023b) and ReClaim. Definitions for Fluency, Correct. and Citation are in Section 4.2.
-4.3 Methods and Baseline
-To evaluate the effectiveness of the proposed method, we tested several models using various generation approaches.
-For each experimental setup, we used the ChatGPT results from the ALCE method as our baseline.
-Additionally, for the two test datasets, we adopted the Oracle-5 paragraphs provided by the ALCE as the reference passages in our model input.
-Prompting
-We directly prompt the model to generate answer text with citations.
-For GPT-4o OpenAI (2023), we employ both zero-shot and three-shot prompting Brown et al.
-(2020), while for GPT-3.5-turbo and open-source models such as Llama3-8B-Instruct Touvron et al.
-(2023) and the vicuna Chiang et al.
-(2023) series, we use 3-shot prompting.
-This approach allows us to comprehensively assess the performance and effectiveness of various models under different prompting scenarios.
-Fine-tuned Model
-[…]
-As outlined in Section 3.4, we fine-tune the same model, generating two separate models: ReferModel and ClaimModel.
-We completed our training on four 80GB A800 GPUs.
-For the ReferModel, we performed full fine-tuning with a learning rate of 2e-5 over 3 epochs.
-For the ClaimModel, we employed Lora tuning Hu et al.
-(2021) with a learning rate of 5e-5 over 5 epochs.
-During the inference phase, we alternated between these two models to interleavingly generate the reference and claim parts.
-To prevent the model from generating overly short or long answers, we set constraints on the number of reference-claim pairs.
-For the ASQA dataset, the model must generate between 2 to 5 pairs.
-For the ELI5 dataset, due to its complexity, the model must generate between 4 to 6 pairs.
-To investigate the roles of the two fine-tuned models, ReferModel and ClaimModel, in the ReClaim w/IG method, we conducted ablation experiments using the Llama3-8B-Instruct model.
-In these experiments, we alternately used only the fine-tuned ReferModel or ClaimModel, while employing the base model with 3-shot prompting to generate the other model’s output.
-For these two fine-tuning methods, we adopted the constrained decoding approach described in Section 3.5 to limit the model’s output in the reference parts.
-Method ASQA ELI5 Average
-Fluency Correct. Citation Fluency Correct. Citation Fluency Correct. Citation Length
-MAUVE EM Rec. CAS CRS MAUVE Claim Rec. CAS CRS MAUVE Claim Rec. CAS CRS
-ALCE 64.4 48.9 74.5 72.7 59.4 21.3 57.8 56.0 61.9 35.1 66.2 64.4 -
-Citation-only 83.5 53.8 65.1 54.6 47.4 17.4 75.1 52.4 65.5 35.6 70.1 53.5 78.9
-Claim-only 43.6 57.7 87.8 72.0 52.7 17.8 81.4 52.2 48.2 37.8 84.6 62.1 121.7
-ReClaim w/IG 86.4 50.5 87.7 77.1 66.8 16.8 84.3 61.8 76.6 33.7 86.0 69.5 86.8
-Table 3: The impact of different components under the ReClaim w/IG method on model generation results.
-Citation-only indicates using the fine-tuned ReferModel to generate references and the base model to generate claims.
-Conversely, Claim-only indicates using the fine-tuned ClaimModel to generate claims and the base model to generate references.
-Method ASQA ELI5
-Length Consistency Attri. Length Consistency Attri.
-Citation Claim CR AR Citation Claim CR AR
-ALCE 536.3 85.5 100.0 91.3 660.0 98.09 100.0 96.9
+ChatGPT 74.9 52.6 72.5 63.4 27.8 17.8 68.6 50.8 30.6 14.1 72.7 55.7
+GPT-4o 91.3 56.6 77.4 58.0 29.7 21.1 70.2 36.8 38.7 18.2 68.0 45.2
 […]
 […]
-[…]
-[…]
-[…]
-[…]
+Llama3-8B 88.1 53.5 92.1 86.1 71.6 17.8 89.9 67.5 63.5 14.0 90.1 68.6
+Table 2: Results on ASQA, ELI5, EXPERTQA. Definitions for Fluency, Correct. and Citation are in Section 4.2.
+4.3 Baselines
+We use only methods that restrict citation sources to the top-k reference passages as our comparison baselines.
+We exclude methods that introduce additional retrieval steps, such as VTG Sun et al.
+(2023) and LLatrieval Li et al.
+(2024), from our comparison.
+Prompting-Based
+Following the ALCE Gao et al.
+(2023b) method, we prompt ChatGPT and Llama3.1-8B-Instruct with few-shot demonstrations that consist of a query, the top-5 retrieved passages, and an answer with inline citations.
+Additionally, we perform four samples of the generation from Llama3.1-8B-Instruct and select the best answer from each case as the experimental result for the w/Rerank method.
+Post-hoc
+Following the ALCE method, we generate answer using the Llama3-8B-Instruct model with the closed-book approach based on the given query.
+Then, for each statement in the answer, we search the top-100 passages for citations.
+We also provide the LLMs with the top-5 reference passages, allowing it to generate the answer based on them and then re-identify citations within the top-5 passages.
+We refer to this method as open-book.
+Training-Based
+Following Self-RAG Asai et al.
+(2023), the LLM is trained to retrieve passages on demand, verify relevance, and generate answers based on the retrieved content.
+We directly provide the selfrag-7B (fine-tuned Llama2-7B model) with the top-5 passages for generation.
+Following the method in Huang et al.
+(2024a), we use the LLM trained with fine-grained rewards (RS+RL, fine-tuned Llama2-7B model) to generate answer with citations.
+Additionally, we compare the fine-grained attribution method FRONT Huang et al.
+(2024b), which is fine-tuned based on the Llama2-7B model.
+4.4 Methods
+We evaluate ReClaim by testing several LLMs with different generation approaches.
+ReClaim prompting
+We directly prompt LLMs to generate answer with citations, using GPT models OpenAI (2023, 2022) and the Llama series Touvron et al. (2023) to evaluate the performance and effectiveness.
+ReClaim ${}_{\texttt{Unified}}$
+We follow the methodology in Section 3.3 and use the training dataset constructed in Section 3.2 to fully fine-tune the Llama3-8B-Instruct model, then conduct one-step generation with citations to evaluate the effectiveness of the ReClaim ${}_{\texttt{Unified}}$ method.
+ReClaim w/IG
+As outlined in Section 3.4, we fine-tune the same base model (Llama2-7B-hf and Llama3-8B-Instruct), generating two separate LLMs.
+During the inference phase, we alternate between these two LLMs to interleavingly generate the reference and claim parts.
+For these two fine-tuning methods, we adopt the constrained decoding described in Section 3.5 to limit the generation of reference parts. For more experimental details, see Appendix C.
+Method ASQA ELI5 EXPERTQA
+Length Consistency Attri. Length Consistency Attri. Length Consistency Attri.
+Citation Claim CR AR Citation Claim CR AR Citation Claim CR AR
+ALCE 536.3 85.5 100.0 91.3 660.0 98.09 100.0 96.9 627.5 115.1 100.0 84.1
+RS+RL 327.0 39.9 100.0 94.5 476.6 80.8 100.0 93.2 501.9 81.9 100.0 95.1
+ReClaim ${}_{\texttt{3-shot}}$ 106.8 59.8 75.5 100.0 162.7 82.1 72.1 100.0 198.7 82.1 77.5 100.0
+ReClaim ${}_{\texttt{Unified}}$ 77.9 52.9 100.0 100.0 139.8 93.1 100.0 100.0 150.9 99.2 100.0 100.0
+ReClaim w/IG 82.8 68.9 100.0 100.0 145.0 104.7 100.0 100.0 157.5 109.1 100.0 100.0
 Table 3: The generated text length, consistency of references, and proportion of attributed answer sentences in different methods.
-CD indicates the use of constrained decoding, while UD indicates the absence of constrained decoding.
+ReClaim ${}_{\texttt{3-shot}}$ does not use decoding constraints, while ReClaim ${}_{\texttt{Unified}}$ and ReClaim w/IG use decoding constraints.
+Method ASQA ELI5 EXPERTQA
+MAUVE EM Rec. CAS CRS MAUVE Claim Rec. CAS CRS MAUVE Claim Rec. CAS CRS
+ReferModel-Only w/Extend 29.4 29.8 70.5 58.9 61.4 13.4 59.9 40.6 50.0 10.1 61.7 44.4
+ReferModel-Only w/Sum 29.2 33.0 94.5 80.8 46.8 14.6 96.2 61.4 49.5 11.0 97.7 69.1
+ClaimModel-Only 43.6 57.7 89.6 77.2 54.5 18.9 84.4 55.1 43.2 14.8 85.6 57.2
+ReClaim w/IG 88.1 53.5 92.1 86.1 71.6 17.8 89.9 67.5 63.5 14.0 90.1 68.6
+Table 4: Ablation study results. The underline indicates the second largest value.
+4.5 Ablation Study
+We primarily conduct ablation experiments on the ReClaim w/IG method to investigate the necessity of training data filtering and fine-tuning the ReferModel and ClaimModel.
+All fine-tuned models are based on the Llama3-8B-Instruct model.
+Pre-Filtered Training Data
+To investigate the necessity of using an NLI model for training data filtering, we fine-tune the LLM on the pre-filtered data while keeping the dataset size and training parameters unchanged.
+The experimental results are shown in Figure 3.
+Figure 3: Comparison of the performance of LLMs trained on the training dataset before and after filtering on the ELI5 dataset.
+The comparative experimental results of the ASQA and EXPERTQA datasets are presented in Appendix Figure 9 and Figure 10.
+The experimental results show that fine-tuning the LLM with data filtered by the NLI model improves its performance across various evaluation metrics.
+Fine-tuned ReferModel Only
+We use the fine-tuned ReferModel to generate reference parts and the base model to generate claim parts.
+When generating claims, it is essential to ensure that it adheres to the information in the preceding reference and maintains fluency with the previously generated claims.
+We explore two prompt strategies:
+  1. 1.
+    Extension: We provide the LLM with the previous reference and all preceding claims, requiring the LLM to extend the claims based on the information in the reference.
+  2. 2.
+    Summary: We provide the LLM with a reference so that it can directly generate a brief summary based on the previous reference.
+Fine-tuned ClaimModel Only
+We employ the 3-shot prompting approach to generate the reference section using the base model and utilize the fine-tuned ClaimModel for generating the claim section.
+The experimental results are shown in Table 4.
+4.6 Faithfulness Analysis
+To further validate our approach in improving the credibility of model responses, we use GPT-4o-mini OpenAI (2023) as our evaluation model to evaluate the faithfulness metric of model answers against complete reference passages using the assessment method proposed by RAGAS Es et al.
+(2023).
+This helps us determine whether the LLM’s generated answers are fully based on the information in the given reference paragraphs.
+We prioritize the CAS metric and use faithfulness as a supplementary measure to evaluate credibility improvements.
+Since faithfulness to a citation implies faithfulness to the original text, a high faithfulness score does not guarantee strong CAS performance, as an answer true to the full document may not align with the selected citation.
+We evaluate the faithfulness metric on the test datasets, and the results for the ELI5 dataset are shown in Figure 4. Results for other datasets are shown in Appendix Figure 7 and 8.
+Figure 4: The x-axis represents the accuracy of the LLM’s responses, while the y-axis shows the faithfulness score.
+For the Self-RAG and RS+RL methods, we use the fine-tuned 7B model, whereas for other methods, we employ Llama3-8B-Instruction as the base model.
+Our method achieved the highest faithfulness score.
+This demonstrates that while our approach may slightly reduce answer quality, it significantly enhances answer faithfulness and minimizes unnecessary hallucinations linked to the LLM’s internal parameters.
+The results show an inverse relationship between the accuracy of LLM-generated answers and their faithfulness to reference passages.
+Higher faithfulness and more granular citations have narrowed the scope of our answer sources, which may contribute to the lower accuracy in LFQA task.
 5 Experiment Results
@@
 5.1 How to Improve the Quality of Answers and Citations?
-The overall performance is presented in Table 2:
-Prompting Works
-Experimental results show that direct prompt models can achieve satisfactory outcomes.
-Compared to the ALCE method using Oracle-5 reference passages, our approach often matches or improves answer flunecy, correctness and citation correctness (CAS).
-In all our prompting experiments, GPT-4o achieved the highest average correctness score (38.9) under 3-shot prompting.
-Additionally, Llama3-8B-Instruct achieved satisfactory results.
-Although it did not surpass the GPT model in terms of average correctness score, it demonstrated the best performance in fluency and CAS.
-Compared to the ALCE method, our model maintained response correctness while improving average fluency and CAS by 22.3% and 17.8%, respectively.
-Although our method performs worse in CRS, the finer granularity of our citations minimizes the impact of redundant content.
-In contrast to paragraph-level citations, where each redundant citation introduces an entire paragraph of irrelevant content, our redundant citations only introduce a single sentence of irrelevant content.
-This does not significantly increase the fact-checking cost.
-Moreover, in many cases, redundant but continuous text helps in locating the citation in the original reference passages and aids in its understanding.
-[…]
-[…]
-[…]
+The overall performance is presented in Table 2.
+ReClaim prompting Works
+Experimental results show that directly prompting LLMs yields satisfactory outcomes.
+Our approach improves average answer fluency and citation accuracy (CAS) compared to baseline methods.
+Notably, Llama3-8B-Instruct surpasses other baselines in CAS scores, including ALCE+ChatGPT.
+Although our method performs worse in CRS, the finer granularity of our citations reduces the impact of redundant content.
+Redundant citations only add a single irrelevant sentence, which does not significantly increase the cost of fact-checking.
+ReClaim ${}_{\texttt{Unified}}$ Cannot Improve ATG
+Experimental results show that the ReClaim ${}_{\texttt{Unified}}$ method significantly reduces citation quality (CAS).
+This indicates that fine-tuning the LLM in this way fails to teach it how to generate claims based solely on the information from the previous reference.
 ReClaim w/IG Improves Attribution
 Experimental results indicate that the ReClaim w/IG method outperforms other methods in two citation quality metrics while maintaining high fluency and correctness scores.
-Specifically, when using the Llama3-8B-Instruct model, our method outperforms the ALCE method with ChatGPT by 23.7%, 30.0%, and 7.9% in fluency, CAS, and CRS metrics, respectively, and only shows a 4.0% decrease in correctness.
-[…]
-As shown in Table 3, the experimental results demonstrate that the ReClaim w/IG method achieves the best performance in metrics of fluency and citation quality when compared to fine-tuning only the RefModel or the ClaimModel methods.
-This indicates the effectiveness of our fine-tuned dual-model approach.
-While the Claim-only method, which fine-tunes only the ClaimModel, yields the highest average answer correctness score (37.8), it comes at the cost of increased response length and reduced fluency.
-Additionally, numerous experiments indicate that improved citation quality often leads to a slight decrease in answer correctness.
-This is expected, as better attribution tends to constrain the source of information for model responses to the selected citations.
+Compared to ALCE using ChatGPT, our method (using Llama3-8B) shows an average improvement of 31.3% on the CAS, 16.7% on the CRS, and 25.7% on the MAUVE across three test datasets, with only a 6.0% decrease in answer accuracy.
+Specifically, we achieved an average CAS score of 90.7 across three test datasets, which is a crucial metric for assessing the degree of text attribution.
+Compared to the ReClaim ${}_{\texttt{Unified}}$ method, the ReClaim w/IG approach’s biggest difference lies in the training and inference strategies during the claim generation phase.
+It filters out extraneous contextual information and trains the LLM to generate claims based solely on the preceding reference part.
+The significant improvements in citation quality demonstrate the effectiveness of the strategy adopted by the ReClaim w/IG method.
+As shown in Table 4, the results of the ablation experiments indicate that fine-tuning two LLMs for alternating generation of references and claims achieves the most balanced performance.
+While ReferModel-Only w/Sum method yields a high citation quality score, it compromises the accuracy and fluency of the answers.
+On the other hand, using ClaimModel-Only method for generation achieves higher accuracy scores, but it negatively affects answer fluency, and the generation often tends to produce excessively long answers.
+Additionally, to ensure that the base model adheres to our format when generating reference parts, we need to provide multiple examples, which reduces the effective context window length.
 5.2 Can ReClaim Enhance the Verifiability and Credibility of RAG-based Question Answering?
-To evaluate the verifiability of model responses and the impact of decoding constraints on their credibility, we measured the average length of citations, the consistency of citation texts (CR), and the proportion of sentences with attribution (AR) under different methods when using constrained decoding (CD) and unconstrained decoding (UD).
-The experimental results are shown in Table 4.
+The results of the average length of citations, citation consistency (CR), and attribution ratio (AR) are shown in Table 3.
 ReClaim Improves Verifiability
-Experimental results indicate that ReClaim’s average citations length is only about 20% of the length produced by the ALCE method, significantly reducing the fact-checking time cost.
-Compared to the ALCE method, our method ensures specific attribution for each response sentence, further improving the verifiability of model’s answer.
-Decoding Constraints Improve Credibility
-[…]
+Experimental results show that ReClaim’s average citation length is about 22% of the ALCE method, significantly reducing fact-checking time.
+ReClaim ensures that each response sentence is supported by a specific citation source.
+By employing constrained decoding, our method ensures that every citation is directly extracted from the original text, significantly reducing the occurrence of hallucinations during the citation generation process.
+Together, these measures collectively enhance the verifiability of the generated answers.
+ReClaim Improve Credibility
+As mentioned in Section 5.1 regarding the improvement of attribution, this enhancement boosts the credibility of the LLM’s responses by providing a definitive source of evidence for the generated answers.
+In addition, as shown in Figure 4, our method increases the faithfulness of the LLM’s responses to the overall reference passages by confining the sources of response information within the chosen citations.
+This improvement reduces the influence of the LLM’s internal knowledge on answer generation, thereby enhancing the credibility of our approach.
+Our method strictly generates answers based on reference passages, aiming to improve citation quality in LLM responses through stringent constraints.
+The accuracy of our answers depends heavily on the information density of the reference passages and cannot rely on the LLM’s internal knowledge, which may also explain the slight decrease in answer accuracy in our approach.
 6 Conclusion
-We propose a attributed text generation method, ReClaim, which adds sentence-level fine-grained citations to model-generated answer in RAG systems.
-This approach alternates between generating citations and answer sentences through Prompting or by fine-tuning the model.
-The results show that our method improves citation quality while maintaining answer quality compared to the baseline method.
-Additionally, our approach significantly reduces the length of citations, thus decreasing the time cost required for fact-checking and further enhancing the verifiability of the model’s responses.
+We propose a attributed text generation method, ReClaim, which adds sentence-level fine-grained citations to LLM-generated answer in RAG systems.
+This approach alternates between generating citations and answer sentences, either through prompting or by fine-tuning LLMs.
+The results show that our method improves citation quality while maintaining answer quality compared to the baseline methods.
+Additionally, our approach significantly reduces the length of citations, thus decreasing the time cost required for fact-checking and further enhancing the verifiability of the LLM’s responses.
 Moreover, by using constrained decoding during citation generation, we ensure that each citation is composed of exact sentences from the source passages.
-Although this paper focuses on long-form question answering task, we hope that our method can be extended to other tasks such as short-form question answering and Multi-document Summarization to improve the verifiability and credibility of model-generated responses.
 Limitations
 In this paper, our training dataset was exclusively targeted at long-form question-answering task, which reduces the generalization ability of our fine-tuning methods.
-Additionally, the construction of our training dataset did not account for the influence of irrelevant passages in the context on the model’s ability to find citations.
-This may weaken the model’s capability to filter out irrelevant information.
-While our approach allows the model to synthesize information from multiple reference sentences for attribution, we did not specifically enhance this capability in our training data construction process and the model’s inference process.
+Additionally, our method requires explicitly outputting the cited sentences, which often leads to generating answers that are more than double the length.
+This results in increased output time for the LLM.
+On one hand, while our approach allows the LLM to synthesize information from multiple reference sentences for attribution, we did not specifically enhance this capability during the training data construction and LLM inference processes.
+On the other hand, our answer template has certain limitations, as not every response generated by the LLM requires a citation from the reference passages.
+Therefore, our method has limitations in scenarios that require synthesizing information from multiple sources or necessitate multi-hop reasoning to draw conclusions.
 Ethics Statement
@@
 We solely utilize the constructed dataset for model training.
-Although paragraph retrieval is not the focus of our work, information retrieved from large corpora may introduce some noise and bias into the responses generated by LLMs.
+Although paragraph retrieval is not the focus of our work, the retrieved information from large corpora may introduce noise and bias into LLM-generated responses.
 To address these issues, we will optimize the data construction process and explore methods for retrieving noise-free and unbiased information.
@@
 Claim-only Use the base model with 3-shot prompting to generate references and the fine-tuned model to generate claims.
-UD Use of constrained decoding.
-CD Absence of constrained decoding.
 Metrics
 MAUVE Measuring the gap between neural text and human text using divergence frontiers.
-EM Rec. Exact match recall rate of gold short answers in the text generated by the model.
-Claim Rec. Recall rate of generated claims in the text generated by the model.
+EM Rec. Exact match recall rate of gold short answers in the text generated by the LLM.
+Claim Rec. Recall rate of generated claims in the text generated by the LLM.
 CAS Correct attribution score, the proportion of sentences predicted as correctly attributed among all the sentences in the answer.
@@
 AR Attribution ratio, the proportion of sentences in the output that are attributed.
-Table 5: A notation table.
-In Table 5, we list the notations and abbreviations in this paper, along with their definitions.
+Table 5: The notation table.
+In Table 5, we list the notations and abbreviations in this paper, together with their definitions.
 Appendix B Details of Training Dataset Construction
-B.1 Citation Selection
-During the training data structuring phase, we initially employ ChatGPT to automatically segment the answers within the WebGLM-QA dataset into clauses.
+B.1 Details of Training Dataset
+The statistics of the training dataset is shown in Table 1
+WebGLM-QA
+The WebGLM-QA Liu et al.
+(2023) dataset is a pioneering resource designed to bolster the development and assessment of web-enhanced question answering systems.
+It distinguishes itself by seamlessly integrating web search functionalities into the QA process, empowering systems to tap into the expansive repository of knowledge on the internet.
+Carefully curated to overcome the shortcomings of existing datasets, WebGLM-QA offers a holistic and pragmatic solution for open-domain question answering tasks.
+It consists of 43,579 high-quality data samples for the train split, 1,000 for the validation split, and 400 for the test split.
+Refer to our paper for the data construction details.
+In WebGLM-QA, the provided answers is generated by the model based on the given reference passages; therefore, we directly use it as our gold answers.
+ELI5
+ELI5 Fan et al.
+(2019) dataset is a benchmark in natural language processing designed for long-form question answering tasks, focusing on complex and explanatory questions (see Table LABEL:table:testdatacase).
+It comprises 270K threads from the Reddit forum “Explain Like I’m Five” (ELI5) where an online community provides answers to questions which are comprehensible by five year olds.
+For ALCE Gao et al.
+(2023b), ELI5 questions were paired with passages from Sphere Piktus et al.
+(2021), a filtered version of Common Crawl.
+For the ELI5 dataset, due to the distributional differences between human answers and the retrieved reference passages, we opted to use Llama-3.1-405B-Instruct to generate answers based on the top-5 reference passages as the gold answers.
+B.2 Citation Selection
+During the training data structuring phase, we initially employ Llama-3.1-405B-Instruct to automatically segment the answers within the WebGLM-QA dataset and ELI5 dataset into clauses.
 Subsequently, for each of these clauses, we undertake a search for corresponding sentence-level, fine-grained citations from the provided reference passages.
@@
 Upon obtaining answer clauses and their respective citations, structure the response in a cot format, as illustrated in Table 8.
-B.2 Citation Filtering
+B.3 Citation Filtering
 After identifying the citation sentences corresponding to each answer clause, we need to filter out citation texts that diverge from the original reference passages or lack sufficient information to substantiate the answer clauses, the algorithm we use is as Algorithm 1.
@@
 9:         end for
-10:         $m\leftarrow M(r_{i},c_{i})$
-11:         if $m=0$ then
+10:         $m\leftarrow p(M(r_{i},c_{i})=1)$
+11:         if $m<0.8$ then
 12:              $flag\leftarrow 0$
@@
 18:end for
+Appendix C Experiment Settings
+C.1 Details of Datasets
+The statistic of the test datasets is shown in Table 9.
+In Figure 5 and Figure 6, we employ bge-v2-m3 Li et al.
+(2023); Chen et al.
+(2024) to compute the relevance between queries and reference passages, presenting the statistical information of both training and test dataset in the form of box plots.
+Below, we provide a detailed description of the test datasets.
+ASQA
+ASQA Stelmakh et al.
+(2022) is the first long-form question answering dataset that focuses on ambiguous factoid questions (see Table LABEL:table:testdatacase).
+It contains 4,353 samples for the train split and 948 samples for the dev split.
+For ALCE Gao et al.
+(2023b), ASQA questions were paired with passages from Wikipedia passages (2018-12-20 snapshot) which purportedly contained the answers.
+EXPERTQA
+EXPERTQA Malaviya et al.
+(2024) dataset is a benchmark designed for evaluating the factuality and attribution capabilities of large language models across diverse domains.
+It contains 2177 long-form questions curated by experts from 32 fields, along with LLM-generated answers that have been revised by these experts to ensure factuality and proper sourcing.
+The dataset aims to provide a high-quality resource for developing and assessing AI systems that can deliver accurate and well-referenced information tailored to the needs of domain specialists.
+ELI5
+Descriptions can be found in the section B.1
+C.2 Details of Methods
+Prompting
+In prompting setting, given a question and reference passages, we simply prompt the model to generate the attributed answer in a prescribed format. The prompt we use is as follows:
+ReClaim ${}_{\texttt{Unified}}$
+In ReClaim ${}_{\texttt{Unified}}$ setting, we directly train the model using the constructed training dataset in Appendix B and employ the same instruction as in the prompting setting for one-step generation of attributed text.
+ReClaim w/IG
+In ReClaim w/IG setting, we train two separate models for generating the reference and claim parts of attributed answers.
+Specifically, we train the model using the constructed training dataset in Appendix B as ReferModel for producing reference parts.
+For ClaimModel, which generates claim parts, we construct a new form of training data based on the original dataset.
+The detailed structure of this tailored training data is shown in Table 10.
+In Table 12, we provide an illustration of an interleaving generation instance.
+C.3 Details of Evaluation Metrics
+EM Rec.
+This metric is employed to assess the accuracy of the ASQA dataset.
+It utilizes a list of short answers (typically in the form of words or phrases) provided within the ASQA dataset to perform string matching within the generated long answers generated.
+The proportion of short answers that can be matched relative to the total length of the short answer list is calculated to determine the metric.
+Claim Rec.
+This metric is designed to evaluate the accuracy of the ELI5 and EXPERTQA datasets.
+Since these datasets do not provide short answer lists but only a long-form answer, ALCE employs InstructGPT Ouyang et al.
+(2022) to sample three sub-claims from the standard answer.
+Then use the NLI model to assess the entailment relationship between each sub-claim and the LLM-generated response.
+The final score for this metric is determined by the proportion of sub-claims that are entailed in the LLM’s answer.
+MAUVE
+We use MAUVE (Measuring the gap between neural text and human text using divergence frontiers) to measure the fluency of the generated text.
+MAUVE is a statistical metric that quantifies the similarity between neural-generated text and human-written text by computing the divergence frontiers between them.
+A higher MAUVE score indicates that the generated text is more coherent and natural, closely resembling human language style.
+CAS
+This metric, consistent with the Citation Recall metric in ALCE, is designed to evaluate the proportion of answer clauses that are sufficiently supported by citations.
+Specifically, for each answer clause associated with a citation, a Natural Language Inference (NLI) model is employed to determine whether an entailment relationship exists between the clause and its corresponding citation.
+If such a relationship is identified, it indicates that the citation adequately substantiates the answer clause.
+The final score for this metric is derived from the ratio of answer clauses that exhibit an entailment relationship with their citations.
+CRS
+This metric aligns with the Citation Precision metric in ALCE, designed to assess whether the citations contain redundancy.
+In the ALCE, citations are identified by paragraph numbers, and redundancy is evaluated by sequentially removing paragraphs.
+However, in our approach, to accommodate sentence-level citations, we first segment the citation into sentences using NLTK.
+We then iteratively remove sentences and verify whether the remaining text still maintains an entailment relationship with the answer clause.
+Finally, the proportion of non-redundant citations is used as the CRS score.
+C.4 Training Details
+We trained the model using llama-factory Zheng et al.
+(2024) on four A800 80G GPUs.
+For the Unified approach, we applied full fine-tuning with a learning rate of 3e-5, 3 epochs, and a total batch size of 128.
+For the ReClaim w/IG approach, we used LoRA to train two models: ReferModel with a learning rate of 5e-5, 5 epochs, and a total batch size of 128, and ClaimModel with a learning rate of 3e-5, 2 epochs, and a total batch size of 128.
+During model inference, we can alternate between two LoRA parameters to reduce memory requirements.
+Figure 5: Reranker score distribution between query and reference passages in the training dataset.
+Figure 6: Reranker score distribution between query and reference passages in the test datasets.
+Appendix D Additional Results
+D.1 Default Datasets
+For our test datasets, in addition to providing gold passages or reranked passages as reference passages, we also tested the results of directly using the top-5 retrieved passages, as shown in Table 6.
+Dataset Type Fluency Correct. CAS CRS
+ASQA Oracle 88.1 53.5 92.1 71.6
+Default 88.1 40.0 91.0 84.4
+ELI5 Oracle 71.6 17.8 89.9 67.5
+Default 75.7 6.5 89.4 66.3
+EXPERTQA Rerank 63.5 14.0 90.1 68.6
+Default 62.9 12.2 87.2 65.7
+Table 6: Results of the default test datasets.
+The results reveal that although the accuracy of the LLM’s responses decreases due to varying quality of reference passages provided to the LLMs, it still maintains high CAS and CRS scores.
+This indicates that our method maintains strong attribution performance even with passages retrieved directly.
+D.2 Additional Baselines
+Due to the differences in datasets and evaluation metrics, we include this baseline in the Appendix.
+Learning to Plan and Generate Text with Citations Fierro et al.
+(2024) conceptualizes plans as a sequence of questions that serve as blueprints for content generation and organization.
+Two variants of blueprint models are introduced: an abstractive model, where questions are generated from scratch, and an extractive model, where questions are copied from input passages.
+For the comparative analysis of this method, we utilized the ASQA and ELI5 datasets, with top-5 reference passages retrieved directly via BM25 and GTR. The experimental results are shown in Table 7.
+ASQA ELI5 Average
+C A C A C A
+LongT5 3B (10-psg) +$Blueprint_{A}$ +Attribution 33.8 77.8 5.2 60.9 19.5 69.4
+ ReClaim w/IG (5-psg) 40.0 87.6 6.5 76.1 23.3 81.9
+ ReClaim w/IG (10-psg) 40.8 88.5 8.0 72.9 24.4 80.7
+Table 7: Comparison with the baseline: Learning to Plan and Generate Text with Citations. C denotes accuracy, while A represents the attribution score, which is the F1 score of CAS and CRS.
+The experimental results indicate that our approach outperforms this method in terms of both answer accuracy and attribution scores.
+D.3 Faithfulness Results
+The experimental results of the faithfulness metric for the ASQA and EXPERTQA datasets are shown in Figure 7 and Figure 8.
+Figure 7: The faithfulness analysis results of the ASQA dataset.
+Figure 8: The faithfulness analysis results of the EXPERTQA dataset.
+Figure 9: The comparative experimental results of the ASQA dataset.
+Figure 10: The comparative experimental results of the EXPERTQA dataset.
 Question Why does a dryer sheet help with my staticy shirt?
@@
 Table 8: Example of the training dataset.
-Appendix C Experiment Settings
-C.1 Details of Datasets
-The statistic of the test datasets of ReClaimis shown in Table 7.
-ASQA
-ASQA Stelmakh et al.
-(2022) is the first long-form question answering dataset that focuses on ambiguous factoid questions (see Table 9).
-It contains 4,353 samples for the train split and 948 samples for the dev split.
-For ALCE Gao et al.
-(2023b), ASQA questions were paired with passages from Wikipedia passages (2018-12-20 snapshot) which purportedly contained the answers.
-ELI5
-ELI5 Fan et al.
-(2019) dataset is a benchmark in natural language processing designed for long-form question answering tasks, focusing on complex and explanatory questions (see Table 9).
-It comprises 270K threads from the Reddit forum “Explain Like I’m Five” (ELI5) where an online community provides answers to questions which are comprehensible by five year olds.
-For ALCE Gao et al.
-(2023b), ELI5 questions were paired with passages from Sphere Piktus et al.
-(2021), a filtered version of Common Crawl.
 Dataset Samples Question Type #passages Average Length
 Question Passages Answer
-ASQA 948 Factoid (ambiguous) Wikipedia (21M) 9.0 500.0 71.8
-ELi5 1000 Why/How/What Sphere (899M) 16.5 500.0 121.5
+ASQA 948 Factoid (ambiguous) Wikipedia (21M) 9.0 517.5 71.8
+ELi5 1000 Why/How/What Sphere (899M) 16.5 546.3 121.5
+EXPERTQA 1000 ambiguous/unambiguous Sphere (899M) 19.4 600.0 152.2
 Table 9: Statistics of the test datasets.
@@
 Table 10: Example of the training dataset for claim generation.
-ASQA Question: Who has the highest goals in world football?
-Reference Passages: [1] "Football Player of the Century", by IFFHS International Federation of Football History and Statistics, 1999, "South America Football Player of the Century", by IFFHS International Federation of Football History and Statistics.
-Pelé’s 1281 goals are recognized by FIFA as the highest total achieved by a professional footballer, although the Soccer Statistic Foundation (rssf) recognizes only 767 goals in official mode, occupying the third place after Josef Bican (805) and Romario (772).
-For his part, Maradona has been named the best soccer player in World Cup history both by The Times and FourFourTwo, publication that also rewarded him as the "Best
-[2] for Christmas, but died less than 2 weeks before that, at the age of 88.
-Josef Bican Josef "Pepi" Bican (25 September 1913 —— 12 December 2001) was a Czech-Austrian professional footballer who played as a striker.
-Rec.Sport.Soccer Statistics Foundation (RSSSF) estimates that he scored at least 805 goals in all competitive matches, which would make him the most prolific scorer of all time.
-Having scored a record 600 league goals and at least 1468 goals overall, the International Federation of Football History & Statistics (IFFHS) awarded Bican the "Golden Ball" as the greatest goalscorer of the last century.
-He
-[3] …
-Answer: The players with the highest all-time goals and highest men’s and women’s international football goals differ.
-The player with the highest all-time men’s football goals is Josef Bican, who in 2020 was recognized by FIFA, the international governing body of football, as the record scorer with an estimated 805 goals.
-Christine Sinclair has the highest goals in women’s international football with 187 and is the all-time leader for international goals scored for men or women.
-Cristiano Ronaldo and Ali Daei are currently tied for leading goalscorer in the history of men’s international football with 109.
-ELI5 Question: How are firms like snapchat, uber etc valued so highly while still not making a profit? Do venture capitalists not expect some form of repayment within a number of years?
-Reference Passages: [1] reporting that the Los Angeles-based company is aiming to raise \$500 million at a valuation of \$16 billion to \$19 billion, making it the third most highly valued tech start-up backed by venture capitalists.
-The Chinese handset maker Xiaomi is valued at \$45 billion, while Uber is estimated to be valued at about \$40 billion, according to data from CB Insights.
-Read MoreVC investment hits \$86B thanks to Uber, Xiaomi Snapchat was valued at \$10 billion in August, according to a Dow Jones report.
-Some of its investors from previous rounds include Benchmark, Lightspeed Venture Partners and Kleiner Perkins Caufield
-[2] Ever wondered how highly valued technology giants like Google and Facebook were able to grow so fast and pay their employees so well in such a short amount of time, or how still growing start-ups like Uber are able to lose 1.2 billion US dollars in just the first half of this year alone and still command a valuation upwards of 50 billion US dollars?
-The answer lies with a special category of investment activity known as venture capital.
-Venture capitalists are professional investors who invest in a number of highly scalable high-risk technology ventures hoping to make a multi-fold
-[3] …
-Answer: Yes.
-Did you watch The Social Network?
-They went a while before introducing ads, so they could make money, as they needed to establish their brand and amass users.
-Once you have dedicated users, introducing ads won’t deter most, but if you are still new, having ads will deter a lot.
-The same goes for Uber, it’s not that they aren’t making money, it’s that they are reinvesting a ton of it to make their service better.
-Table 9: Example of the test datasets.
-C.2 Details of Methods
-Prompting
-In prompting setting, given a question and reference passages, we simply prompt the model to generate the attributed answer in a prescribed format. The prompt we use is as follows:
-[…]
-[…]
-ReClaim w/IG
-In ReClaim w/IG setting, we train two separate models for generating the reference and claim parts of attributed answers.
-Specifically, we train the model using the constructed training dataset in Appendix B as ReferModel for producing reference parts.
-For ClaimModel, which generates claim parts, we construct a new form of training data based on the original dataset.
-The detailed structure of this tailored training data is shown in Table 8.
-In Table 10, we provide an illustration of an interleaving generation instance.
+Table 11: Example of the test data.
+Question: When data is compressed or zipped, what is actually happening to the data?
+Reference Passages:
+[1] Title: Lossless vs. Lossy Compression: What’s the Difference?
+Text: data in several different ways, balancing fidelity and efficiency for functional and presentable data on the egress end.
+A common implementation of lossless file-compression includes the use of Huffman coding, whose redundancy-limiting algorithm recognizes patterns in groups in order to conserve time, space and other resources.
+The model is able to compress and decompress digital media such that the output perfectly matches the input.
+The Zip file archival tool is a well-known format that supports lossless compression.
+Zip files compress digital media into much smaller data (often given the 2̆018.zip2̆019 file extension) that can be uncompressed into their original form,
+[2] Title: What is MP3? (Designing Web Audio)
+Text: space, try encoding in mono.
+To make sure you trap all possible spatial data, encode in stereo mode.
+Most users find that joint stereo is adequate for most purposes.
+As mentioned earlier, compressing a WAV file with zip doesn’t shave much off the file size, which is why psychoacoustics are employed.
+However, the MP3 encoding process actually does employ the classic Huffman encoding algorithm.
+After all psychoacoustic methods have been applied, the Huffman encoding pass seeks out and compresses any remaining redundancies in the bit pattern.
+It’s as though zip-type encoding were being run internally on the psychoacoustically encoded data
+[3] Title: US Treasury: addendum | Zip (File Format) | Comma Separated Values
+Text: most large files on the Internet are compressed.
+File compression reduces the size of a file and the time it takes to download.
+Compression software uses complex mathematical equations to scan a file for repeating patterns in the data.
+It replaces the data with smaller codes that take up less room.
+This can dramatically decrease the size of report files.
+Viewing a 2̆01cZipped2̆01d file To view compressed files, you need a compatible decompression program that can read these codes and convert the data back to its original form.
+To decompress a .zip file you need a utility like WinZIP2̆122 (an
+[4] Title: An introduction to data compression
+Text: data and try to find redundant parts of that data and remove those redundancies.
+That wasn’t too simple, but the definition needs to be somewhat complex in order to differentiate compression from some other technologies.
+2.
+How is compression accomplished?
+Can you give us a picture of what actually happens to the data?
+Basically, they’re trying to find common pieces of data blocks that they can get rid of, shrink, remove or substitute with smaller patterns.
+The more of those things they can find, the more it can compress.
+A perfect example to illustrate how that works is if you
+[5] Title: Introduction to Data Compression - Data Compression | Coursera
+Text: we get into the topic.
+so, just specific applications that, maybe are familiar for file compression.
+All file systems and, and disks have built-in, compression technologies.
+Such as, as zip, and b-zip and many others of similar type.
+Technologies.
+And the multimedia, everybody’s familiar with, the JPEG and MP3 and MPEG, and all those sorts of things for images, sound and video.
+Those are all about data compression.
+and for communication, now, that’s, what has, what enabled, fax, and also enables new technologies, like Skype, the ability to, reduce the, amount of data that you actually need to send.
+for (Irrelevant passage)
+Gold Answer: For most algorithms, the compressor will search for a pattern that appears multiple times within the data, and then replace that pattern with a shorter symbol, and create an entry in a "dictionary" recording what each symbol means.
+Compressing algorithms can get a little more complex than that, but that’s the general idea.
+Model Answer: According to the citation: <reference> Compression software uses complex mathematical equations to scan a file for repeating patterns in the data.
+It replaces the data with smaller codes that take up less room.
+</reference> We can know that: <claim> Compression software uses complex mathematical equations to scan a file for repeating patterns in the data and replaces the data with smaller codes that take up less room.
+</claim> According to the citation: <reference> A common implementation of lossless file-compression includes the use of Huffman coding, whose redundancy-limiting algorithm recognizes patterns in groups in order to conserve time, space and other resources.
+</reference> We can know that: <claim> A common implementation of lossless file-compression includes the use of Huffman coding, whose redundancy-limiting algorithm recognizes patterns in groups in order to conserve time, space and other resources.
+</claim> According to the citation: <reference> After all psychoacoustic methods have been applied, the Huffman encoding pass seeks out and compresses any remaining redundancies in the bit pattern.
+</reference> We can know that: <claim> After all psychoacoustic methods have been applied, the Huffman encoding pass seeks out and compresses any remaining redundancies in the bit pattern.
+</claim> According to the citation: <reference> Basically, they’re trying to find common pieces of data blocks that they can get rid of, shrink, remove or substitute with smaller patterns.
+The more of those things they can find, the more it can compress.
+</reference> We can know that: <claim> The more of these common pieces of data blocks that can be found, the more it can compress.
+</claim>
 Step Model Input Output
-1 ReferModel Instrction: Given the Question and References below, provide an answer for the Question that is generated using information exclusively from the References(some may be irrelevant). …
+1 ReferModel Instruction: Given the Question and References below, provide an answer for the Question that is generated using information exclusively from the References(some may be irrelevant). …
 # Question: Who has the highest goals in world football?
@@
 </claim>
-3 ReferModel Instrction: Given the Question and References below, provide an answer for the Question that is generated using information exclusively from the References(some may be irrelevant). …
+3 ReferModel Instruction: Given the Question and References below, provide an answer for the Question that is generated using information exclusively from the References(some may be irrelevant). …
 # Question: Who has the highest goals in world football?